### v1

# Fully Distributed Fog Load Balancing with Multi-Agent Reinforcement Learning

Abstract
Real-time Internet of Things (IoT) applications require real-time support to handle the ever-growing demand for computing resources to process IoT workloads. Fog Computing provides high availability of such resources in a distributed manner. However, these resources must be efficiently managed to distribute unpredictable traffic demands among heterogeneous Fog resources. This paper proposes a fully distributed load-balancing solution with Multi-Agent Reinforcement Learning (MARL) that intelligently distributes IoT workloads to optimize the waiting time while providing fair resource utilization in the Fog network. These agents use transfer learning for life-long self-adaptation to dynamic changes in the environment. By leveraging distributed decision-making, MARL agents effectively minimize the waiting time compared to a single centralized agent solution and other baselines, enhancing end-to-end execution delay. Besides performance gain, a fully distributed solution allows for a global-scale implementation where agents can work independently in small collaboration regions, leveraging nearby local resources. Furthermore, we analyze the impact of a realistic frequency to observe the state of the environment, unlike the unrealistic common assumption in the literature of having observations readily available in real-time for every required action. The findings highlight the trade-off between realism and performance using an interval-based Gossip-based multi-casting protocol against assuming real-time observation availability for every generated workload.
Index Terms:
Internet of Things, Fog Computing, Load Balancing, Reinforcement Learning, Multi-Agent Learning
I Introduction
In the era of interconnected devices, Internet of Things (IoT) has emerged as a pervasive technology, facilitating seamless communication and data exchange among various devices. However, the surge in IoT devices has posed significant challenges in efficiently managing and processing the generated data. Fog computing has emerged as a promising paradigm to address these challenges by distributing computing resources closer to data sources. Unlike the Cloud, Fog nodes can serve time-sensitive and computationally intensive applications of resource-limited IoT devices.
Fog Computing necessitates efficient resource management due to its distributed nature and latency-sensitive applications. Distributing IoT load efficiently between available Fog resources is essential to reduce end-to-end latency and enhance the system performance. High workload generation rates can increase the waiting delay due to the accumulation of tasks in the queues of Fog nodes. Hence, minimizing the waiting delay is the key to reducing the overall end-to-end execution delay of IoT workloads.
Load balancing in Fog computing environments plays a crucial role in ensuring optimal resource utilization and timely processing of IoT data; this allows for minimizing task completion time and improving overall network performance. Traditional load-balancing techniques often struggle with the dynamic nature of Fog environments, characterized by heterogeneous resources, limited communication bandwidth, and fluctuating workloads. This is where Reinforcement Learning (RL) comes into play, as it allows agents to learn optimal decision-making policies through interaction with the environment. However, managing Fog and IoT resources must be done at the network edge, in a distributed manner. This will avoid the need for distant decisions, taking full advantage of the distributed computing capabilities of Fog systems [1].
Unlike centralized RL-based load-balancing solutions, fully distributed Multi-Agent RL (MARL) overcomes scalability limitations by allowing autonomous agents to learn from their local observations and interactions. These agents can be deployed in network Access Points (APs), i.e., Base Stations, to distribute the load their associated end devices generate. This paper presents a novel fully distributed approach for Fog load balancing using MARL. Our approach introduces distributed self-learning agents capable of efficiently distributing IoT load across the Fog environment. These agents incorporate self-adaptation mechanisms, leveraging practical lifelong transfer learning to adapt to dynamic environmental changes seamlessly. This distributed solution outperforms the centralized lifelong learning agent proposed in our previous work [2].
Available MARL solutions in the literature provide collaboration between RL agents through one or more centralized components, like a centralized policy and a centralized experience replay buffer, which is not practical for a global-scale Fog load-balancing solution. Without fully decentralizing decision-making, the communication overhead grows exponentially with the number of agents and Fog nodes in the system, especially when the agents need to communicate with each other. In addition, proposing fully independent agents lowers the state and action spaces of individual agents to reduce the computational complexity of the solution. This is done by considering smaller sets of Fog nodes, e.g., within a certain proximity, as possible candidates for each group of agents (see Fig. 1).
Figure 1: A simplified Fog network with overlapping collaboration regions.
Fig. 1 shows the formation of smaller collaboration regions, where the group of agents accessing the same set of candidate Fog nodes can form a collaboration region. Each region contains one or more agents that need to distribute the load between one or more candidate Fog nodes. In this setup, an agent can only be part of a single collaboration region. Candidate Fog nodes can be shared between two or more overlapping regions, while an isolated collaboration region does not share its set of candidate Fog nodes with other regions. Allowing the agents to independently, i.e., blindly, learn to collaborate in each collaboration region leads to the global optimization of the whole system.
The independent agents must learn to optimize the same objective of minimizing the overall system waiting delay. To avoid unnecessary communication overhead, each agent observes the same state from the environment while independently taking actions and receiving rewards for those actions. Hence, our independent distributed agents can work without sharing their knowledge, experience, or actions, avoiding the need for a centralized coordination entity and inter-agent communication. In addition, a realistic multi-cast approach for observing the environment is proposed in this work, compared to unrealistic real-time observations often considered in the literature.
This paper proposes the first fully distributed, efficient, and self-adapting MARL load distribution solution for realistic global-scale Fog Computing implementations without centralized components or inter-agent communication. Through empirical evaluation, we show the effectiveness of our novel truly distributed MARL-based load-balancing solution and compare it with baseline solutions and a centralized state-of-the-art RL solution. We also show the performance of observing the environment using realistic multi-cast intervals compared to unrealistic real-time observations. Hence, the key contributions of this paper can be summarized as follows:
  • Propose fully distributed self-learning LB RL agents that efficiently distribute IoT load into a Fog Computing environment.
  • Augment the independent distributed agents with lifelong self-adaptation to dynamic environment changes using transfer learning.
  • Provide a global-scale solution by allowing the agents to work independently in small collaboration regions, unlike centralized architectures.
  • Evaluate realistic interval-based collection of observations using Gossip multi-casting protocol compared to the common practice in the literature of unrealistically providing this information in real-time.
The rest of the paper is organized as follows. Section II presents the related work. Section III presents the system model. Section IV presents the proposed approach. Section V presents the evaluation of our proposed approach. Finally, Section VI concludes the paper.
II Related Work
In recent years, researchers have made significant contributions toward developing efficient resource management solutions for Fog Computing environments. One such approach is centralized RL, where a single agent manages the entire network. However, centralization causes scalability limitations and single-point-of-failure vulnerabilities. Using a centralized load-balancing node in large-scale Fog networks increases the time needed to make the load assignment decision, increasing end-to-end execution delay. Mitigating the limitations of single-agent RL solutions requires distributed multi-agent solutions, where multiple agents collaborate to manage Fog resources across the network.
Jain et al. [3] proposed training an agent for each Fog node to mitigate the delay of decision-making in centralized solutions. However, they trained all the agents offline on a centralized cloud data center to later distribute the trained models to each Fog node to make offloading decisions locally with minimal delay. However, training these agents using this approach, and retraining them in case of changes in the environment, requires the collection of local experience trajectories of each agent, i.e., observations, actions, rewards, and next observations. Transferring such information to a centralized entity to retrain these agents is inefficient compared to local retraining, especially for highly dynamic environments that keep changing over time.
Determining the optimal timing for retraining these agents in the centralized entity presents a critical trade-off between decision accuracy and the significant retraining overhead. In addition, an RL decision can only be made when the observation of the current step is available, along with the reward of the previous step. Observing the state, for each decision, of every Fog node in large-scale Fog networks will cause a significant delay, especially with frequently generated workloads [4]. This delay is often not considered in the literature when proposing load-balancing solutions with real-time decision-making. Hence, the effectiveness of these solutions in practice can not be guaranteed without considering the time needed to transfer Fog load provisioning information to all RL agents.
Gao et al. [5] used MARL to address the scalability problem of RL-based load-balancing solutions. They provide task offloading and resource allocation in large-scale heterogeneous Multi-access Edge Computing (MEC) systems. In their design, each edge server is regarded as an actor agent while still considering a centralized critic, where the critic is a neural network used to evaluate the actions of actor networks [6]. This centralized critic facilitates the coordination of offloading policies among those distributed actors. Besides not being fully distributed, the authors assumed that observations are readily available when the agents need to make assignment decisions. This is a common unrealistic assumption in the literature, where the delay of exchanging Fog status information to every agent is often ignored.
Baek et al. [7] proposed a centralized critic network with distributed individual actor networks. Those distributed agents are located in APs and they cooperatively learn to solve the joint link and server load-balancing problem. The policy of the distributed cooperative agents is optimized using parameter-sharing and distributed execution; the agents communicate among themselves while evaluating a common value function. The authors assume that updated policy parameters are shared between the agents by exchanging communication messages over limited bandwidth channels. However, they did not show the transmission delay associated with those communication messages, which augments when scaling up the solution for a larger-scale Fog environment. Such a delay can have a significant impact on the collaboration of the agents due to sharing outdated parameters.
Similarly, Goudarzi et al. [8] proposed a MARL framework to improve the execution delay of heterogeneous Fog services using multiple actors and a centralized critic. They deployed distributed agents across different Fog nodes to work in parallel, i.e., brokers that make service offloading decisions. Then, they train a centralized service offloading model, i.e., the learner of the target policy, using the experience trajectories generated by the distributed actors. But again, they did not consider the delay caused by exchanging such information in a large-scale environment with high request rates. In addition, they did not consider the delay when updating the common value function of each actor from the centralized critic network.
Wei et al. [9] proposed a scheme to improve vehicular Fog resource utilization in large-scale Vehicular Fog Computing (VFC) networks. They used MARL to decide on multi-tier task offloading in a distributed manner. Like most MARL-based solutions in the literature, they used the approach of centralized training with decentralized execution (CTDE). They used the attention mechanism to share the agents’ parameters with a joint critic network. This joint network incorporates the learning process of distributed gated actors that exploit the local computational potentials.
Zhang et al. [10] also used CTDE-based MARL to reduce the long-term average delay and energy consumption while improving the throughput of dynamic Edge networks using online Request Scheduling. Centralized training was adopted to improve the long-term system performance by realizing the implicit cooperation among edge nodes. This cooperation was obtained by transmitting the replay buffer data of each agent into a shared training server to train a common policy. During distributed execution, on the other hand, only local environment information is needed for each edge node to make the request scheduling decision. However, the authors did not discuss the scope of the local information, and the delay associated with transmitting such information over the network. In a large-scale network, this delay is not tolerable for real-time decisions required by real-time applications.
Lu et al. [11] proposed MARL-based dynamic scheduling to handle the massive number of service requests by the ever-increasing number of vehicles in Internet-of-Vehicle (IoV) networks. They leveraged CTDE to learn high-quality scheduling rules for the service function scheduling problem. In their architecture, the generated scheduling actions of the distributed policies are evaluated by separate critic networks deployed in a centralized controller node. Similarly, Cui et al. [12] proposed offline CTDE to cooperatively optimize the offloading strategy for computation-intensive and delay-sensitive IoV applications. Offline CTDE helped reduce the communication between the agents while increasing the stability of the environment. However, offline execution does not allow the agents to dynamically adapt to environment changes. In addition, their solution will not scale to larger networks as all their agents must observe the global state of the environment, which are assumed, unrealistically, readily available for every decision to be made.
CTDE was also used by Gao et al. [13] to predict the queuing delay of edge servers by training a centralized critic in the Cloud. To evaluate each agent’s policy, the critic network needs to collect observations and their corresponding actions from all agents in the system. Then, the agents independently make their decisions using their policies distributively. Likewise, Suzuki et al. [14] used CTDE to send learning samples from each agent to a centralized training entity to avoid the non-stationary problem of independently trained agents. A non-stationary environment is when the new state depends on the joint action of multiple agents instead of a single agent’s behavior. To learn and make offloading decisions, the agents in these approaches observe task information, network link utilization, and Fog resources utilization, which are all unrealistically assumed available in real-time for all agents decisions.
Mangiaracina et al. [15] proposed adaptive MARL to manage data placement among Fog nodes considering Quality of Service (QoS) requirements. Unlike CTDE, they enabled cooperation between multiple agents using the concept of a shared repository accessible by all agents. This repository contains knowledge bases and event logs shared between those agents to improve their decision-making. The authors claim that this centralized repository does not affect the overall performance as a bottleneck in the architecture. However, they did not consider the time each agent needs to access information from this repository.
To coordinate the decisions of independent agents, Yang et al. [16] proposed that the decisions of each agent must depend on shared global state information. They proposed a centralized migration scheduling model to collect server load information and calculate the migration likelihood. They estimate the global information in a centralized model to reduce the communication overhead caused by broadcasting this information from each agent to the centralized critic network. They showed a 20% gap in performance between their approach and the optimal solution but with only half the decision-making time, where the optimal solution is a centralized decision-making method having the states and actions available for all agents.
Most existing MARL solutions adopt a centralized training approach, no fine-tuning for dynamic adaptation, and unrealistic assumptions for observation availability (see Table I). The centralized component in these solutions raises several challenges, especially in large-scale environments, including computational complexity, exploration of larger state and action spaces, and the communication overhead of inter-agent exchanges. Despite the efforts to mitigate these challenges in CTDE-based solutions, a truly distributed solution remains preferable to enhance scalability and decision-making time. With true distribution of decision-making, each agent makes decisions independently, without having access to the state of other collaborating agents in the system.
TABLE I: Comparison of state-of-the-art MARL-based Fog load balancing solutions.
SOTA Training Retraining Information Sharing Observation Availability
Jain et al. [3] Centralized No A shared trained model Unrealistically instant
Gao et al. [5] Centralized No Parameter Sharing (Centralized) Unrealistically instant
Baek et al. [7] Centralized No Parameter Sharing (Inter-agent) Unrealistically instant
Goudarzi et al. [8] Centralized No Experience Sharing (Centralized) Unrealistically instant
Wei et al. [9] Centralized No Parameter Sharing (Centralized) Unrealistically instant
Zhang et al. [10] Centralized No Experience Sharing (Centralized) Unrealistically instant
Lu et al. [11] Centralized No Experience Sharing (Centralized) Unrealistically instant
Cui et al. [12] Centralized No Experience Sharing (Centralized) Unrealistically instant
Gao et al. [13] Centralized No Experience Sharing (Centralized) Unrealistically instant
Suzuki et al. [14] Centralized No Experience Sharing (Centralized) Unrealistically instant
Mangiaracina et al. [15] Distributed No Shared knowledge base Unrealistically instant
Yang et al. [16] Centralized No A shared trained model Unrealistically instant
Our fully distributed approach Distributed Lifelong learning None Interval-based observations
This paper proposes a fully distributed solution that efficiently coordinates decision-making among independent agents without reliance on centralized components or inter-agent communication. This truly distributed MARL framework effectively handles the complexities of Fog environments, scales well to a global-scale solution, and minimizes communication overhead and decision latency. In addition, this work presents the first attempt to evaluate a realistic way to observe the environment and the reward of the actions of each agent based on a multi-cast Gossip protocol.
III The Fog System Model
This section defines the system model of the Fog network and the formulation of the load balancing problem. Table II shows the definitions of the mathematical symbols used throughout this paper. We reproduce the realistic Fog environments used in our previous works [17, 18] to compare the proposed fully distributed multi-agent solution with the centralized single-agent solution in [17].
TABLE II: List of mathematical notations.
Symbol Definition
$N$ Set of AP, Fog, and Cloud nodes ($N^{AP},N^{F},N^{C}$)
$IPT_{n}$ Compute capabilities of node $n$ (Instructions/timestep)
$RAM_{n}$ Memory capabilities of node $n$
$L$ Set of Network links connecting the nodes
$BW_{l}$ Bandwidth of link $l$
$PR_{l}$ Propagation delay of link $l$
$DA$ Set of distributed applications
$M_{da}$ Set of modules of application $da$
$D_{da}$ Set of dependencies between the modules of $da$
$w$ A category from the set of workload categories $W$
$\beta$ The scale parameter of the exponential distribution
$\mathbb{Q}_{n}$ The number of queued jobs in node $n$
$s_{t}$ The observed state at step $t$, from a set of possible states $S$
$a_{t}$ The agent’s action at step $t$, from a set of possible actions $A$
$r_{t}$ The agent’s immediate reward at step $t$
$G$ The discounted sum of immediate rewards (return)
$\gamma$ The discount factor of the return
$\mathcal{F}$ Set of candidate Fog nodes in a collaboration region $\mathcal{R}$
$\mathcal{F}_{S}$ Shared Fog nodes between two collaboration regions
$\pi(s,a)$ The agent policy for taking action $a$ in state $s$
$Q_{\pi}(s,a)$ The value for using $\pi$ after taking action $a$ in state $s$
$Q^{*}$ The optimal $Q$-Value for using the optimal policy $\pi^{*}$
$\phi(s,a)$ A mapping of the state-action pairs
$\theta$ The neural network weights for the state-action pairs
$\alpha$ The learning rate for updating the $Q$-Value
$\max_{a}Q$ The maximum $Q$-Value in state $s$ for all actions
$\operatorname*{arg\,max}_{a}Q$ The action that maximizes the $Q$-Value in state $s$
$E_{\mathcal{T}}$ The new environment (target task $\mathcal{T}$)
$\mathcal{I}_{\mathcal{S}}$ Information from the source task $\mathcal{S}$ (old environment)
$\mathcal{B}$ Experiences stored in the agent’s replay buffer
The Fog environment has a mesh-like topology, with a central cloud, interconnected Fog nodes, and APs that work as gateways for their associated IoT devices. Load-balancing agents are deployed in these APs, where each agent makes its own decisions to distribute the load of its associated IoT devices between the Fog nodes in its collaboration region, i.e., its candidate Fog nodes. Agents in each collaboration region share a set of candidate Fog nodes, requiring the agents to optimize the resources of all candidate Fog nodes collaboratively. To be fully distributed, the agents must collaborate blindly, i.e., without sharing their experience or decision-making policies.
The Fog network is defined by the set of $N=n_{1},n_{2},\cdots,n_{z}$ nodes described by their compute $IPT_{n}$ and memory $RAM_{n}$ resources [17, 18]. The Nodes, including $N^{AP}$ AP nodes, $N^{F}$ Fog nodes, and $N^{C}$ Cloud nodes, are connected through $L=l_{1},l_{2},\cdots,l_{z}$ bidirectional links. Each link is characterized by the pair ($n_{i}$, $n_{j}$) that it connects, its bandwidth $BW_{l}$, and its propagation delay $PR_{l}$. The network has a set of distributed applications $DA=da_{1},da_{2},\cdots,da_{z}$ that simultaneously run in the system; each application has a set of modules $M_{da}$ and dependency messages $D_{da}$ between these modules, i.e., workloads or requests.
Nodes and network links have heterogeneous resources and must serve three different workload categories $|W|$ each with different compute requirements, i.e., light, medium, and heavy workloads. Workloads are generated as a Poisson Point Process using exponential distributions with three different scale parameters $\beta$ to simulate high, medium, and low workload generation rates, allowing for dynamic representation of job arrivals in our Fog environment. Nodes and links are modeled with M/M/1 queuing models to handle the generated workloads, giving the ability to provide insights into performance metrics such as queue lengths and waiting times. The agents must learn to minimize the number of queued jobs $\mathbb{Q}_{n}$ in the queue of each Fog node $n$ to minimize the workload waiting delay, reducing the end-to-end execution delay over the system [17].
To optimize this objective, each agent learns a policy to select the best Fog node for each generated workload. To do this, each agent in a collaboration region observes the number of jobs queued in every candidate Fog node in that region. These observations represent the state of the environment at each step, and it is the same for all agents in a collaboration region. This state is used to select an action $a$ from the set of possible actions $A$, corresponding to the set of candidate Fog nodes $N^{F}$ in that region. Each agent learns to select the action that maximizes its performance measure, i.e., the reward, which is evaluated at the end of every decision step. For each agent, the reward is the number of jobs queued in the Fog node selected in its previous action. To reduce the average system waiting delay, this number is negated to be maximized by the RL algorithm. This reduces the accumulation of workloads in Fog nodes, lowering overload probability and overall execution delay.
IV Distributed RL Load Balancing
A fully independent Double Deep Q-Learning (DDQL) agent is deployed in each of the $N^{AP}$ APs of the Fog network. The state, action, and reward function of every agent at every decision step $t$ are defined as follows:
  • State: $s_{t}=(w_{t},\mathbb{Q}_{n,t}:\forall n\in N^{CF})$, where $w$ represents the category of the workload to be assigned and $\mathbb{Q}_{n}$ represents the number of queued jobs in candidate Fog node $n$.
  • Action: $a_{t}\in A$ is the decision to assign the current workload to the best candidate Fog node from the set of candidates considered for that agent.
  • Reward: $r_{t}=-\mathbb{Q}_{a_{t-1},t}$ represents the number of queued jobs in the Fog node previously selected by that agent, i.e, the reward at decision step $t$ is the performance measure of the action taken at decision step $t-1$ (it is part of the observed state, and thus, does not need to be transmitted separately to each agent). The negation is used to represent the reward as a cost to be minimized instead of a utility to be maximized. The accumulation of these rewards is the expected discounted return $G=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, where $\gamma$ defines the importance of future rewards compared to recent ones.
In centralized single-agent solutions [2, 17], the state space consists of the number of queued jobs from all Fog nodes at each decision step, i.e., $\mathbb{Q}_{n,t}:\forall n\in N^{F}$. The centralized agent also requires the identity of the AP connected to the source IoT device, i.e., workload from $n^{AP}_{t}$ to be assigned at decision step $t$. When shifting to a distributed multi-agent solution in large-scale networks, considering all Fog nodes for every agent can unnecessarily increase the state and action spaces of each agent. However, it is sufficient to design the agents with feasible subsets of Fog nodes within a certain proximity. This leads to creating multiple groups of agents, each observing the queues of their own set of candidate Fog nodes $\mathbb{Q}_{n,t}:\forall n\in N^{CF}$.
In addition to increasing the state and action spaces, a centralized agent causes a significant delay for each workload assignment decision due to the three-way communication between the centralized agent and the workload AP (see Fig. 2LABEL:sub@fig:Topology_WF_C). For every generated workload in centralized load distribution solution, an assignment request is sent to the centralized agent along with the workload category and the identifier of its AP. After receiving this request, the centralized agent predicts the best Fog node for that workload using the provided information and the observed number of queued jobs in each Fog node in the system. The agent then responds to the requesting AP with the identifier of the selected Fog node. Finally, the AP assigns and sends the workload to the selected Fog node and waits for the next generated workload.
(a) Centralized agent
(b) Distributed agents
Figure 2: Workflow of workload distribution.
With a distributed load balancing solution, however, each agent has this prediction capability locally, and hence, each agent makes its own assignment decisions (see Fig. 2LABEL:sub@fig:Topology_WF_D). As soon as the AP receives a workload, it predicts the best candidate Fog node for that workload by providing the workload category $w$ and the number of queued jobs in each candidate Fog node. The workload is then immediately assigned and transmitted to the selected Fog node, minimizing the decision delay compared to a centralized solution.
In addition to the decision-making delay induced by the agent’s location, i.e., centralized vs. distributed, observing the number of queued jobs in each Fog node causes a significant delay, often neglected in the literature. In a centralized solution, each Fog node sends its queue information to the centralized agent as in Fig. 3LABEL:sub@fig:Topology_info_C. In a distributed solution, however, each agent/AP needs to receive this information from all candidate Fog nodes (see Fig. 3LABEL:sub@fig:Topology_info_D). To do this in a real-world environment, Fog nodes need to multi-cast their queue information, introducing a significant delay compared to a centralized solution.
(a) Centralized agent
(b) Distributed agents
Figure 3: Transmission of Fog queue information.
Frequently multi-casting the observations, i.e., for every generated workload, in large-scale networks consumes a significant amount of network bandwidth. Hence, providing this information in real-time for every workload assignment decision becomes infeasible, especially with a large number of Fog nodes and IoT devices, and high workload generation rates. Therefore, addressing the frequency of transferring this information between the agents becomes essential. In addition, dividing large-scale networks into smaller subsets of collaboration regions limits the transfer of this information to fewer agents within a smaller geographical scope, thereby reducing the communication overhead.
A distributed load balancing solution allows the network to be clustered into smaller, possibly overlapping, regions based on geographical proximity between Fog nodes and APs. This significantly reduces the state and action spaces by only considering feasible Fog nodes for the agents in each collaboration region, which helps reduce the computational complexity of the solution. The agents in each region will then collaborate within their respective boundaries to optimize the resources of their feasible candidate Fog nodes. Unlike centralized load balancing solutions, where all Fog nodes must be included in the observation of the centralized agent, whether they are feasible options for all APs or not.
In this paper, we evaluate our fully distributed solution on a single collaboration region, where all agents collaborate to optimize the resources of the candidate Fog nodes in their region. To show the scalability of this solution to a larger network with multiple collaboration regions, these agents must be able to independently collaborate to minimize the number of queued jobs in each candidate Fog node. To learn independently, the agents must collaborate blindly, i.e., without knowing the policies and actions of the other agents in the system.
Lemma IV.1.
Given two non-overlapping sets of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\emptyset$ in two collaboration regions. Each region is managed by a group of agents that blindly collaborate to optimize the resources of their candidate Fog nodes without knowing the actions of other agents in the system (see Fig. 4). If the two regions share a subset of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\mathcal{F}_{S}$, the agents of each region can still blindly collaborate to optimize the resources of their set of candidate Fog nodes and the shared set of candidate Fog nodes.
Proof.
Assuming the agents in an isolated region can efficiently manage the resources of their candidate Fog nodes by blindly collaborating without the need to observe the actions, nor the mechanism of making those actions, of other agents in their collaboration group. This assumption holds for any number of isolated regions each with their own set of candidate Fog nodes. Now, assume that a region $\mathcal{R}_{1}$ with $\mathcal{F}_{1}$ candidate Fog nodes share $\mathcal{F}_{S}$ candidate Fog nodes with region $\mathcal{R}_{2}$ (see Fig. 4). Suppose there exists an agent in $\mathcal{R}_{2}$ that does not collaborate with the agents in its group leading to inefficient management of the Fog resources in $\mathcal{F}_{2}$, including those in $\mathcal{F}_{S}$. This implies that there is a breakdown in the collaborative optimization process. However, this contradicts our initial assumption that within each region, the agents efficiently collaborate to optimize the resources of their candidate Fog nodes. Therefore, we conclude that the agents of both regions will collaborate to optimize the shared resources in $\mathcal{F}_{S}$ as well as those of their own. Therefore, the lemma holds. ∎
Figure 4: Overlapping vs. non-overlapping sets of candidate Fog nodes for two coordinating regions.
In this work, we provide empirical evidence under a realistic setup supporting the ability of our independent distributed RL agents to optimally distribute the load within a collaboration region. Each agent acts according to the current environment state, and the learned policy $\pi$. In real-world environments, it is impractical to observe the state of every candidate Fog node for every workload assignment decision due to the network delay, which is intolerable for real-time applications. Observing the state of all Fog nodes in large-scale networks necessitates waiting for the state of the last node to transmit its load information to complete the environment observation. Hence, the effect of the observation delay is impacted by the scale of the Fog system, the number of IoT devices, and the rate of workload generation. In the literature, however, observations are always assumed to be immediately available for the agent at each decision step [19], rendering those solutions impractical for real-world implementations.
To provide a practical solution with real-world constraints, we use a realistic approach to observe the number of jobs queued in each Fog node. In this approach, the agents collect observations using interval-based multi-casting protocols, i.e., sending queue information on predetermined intervals. Using the same observation for multiple decision steps will lead to sub-optimal solutions, depending on the frequency of collecting the observations. Hence, the smallest practical interval for the multi-casting protocol must be considered to transfer the queue information to all agents. In this work, for instance, Fog nodes transfer their queue information every 3 seconds, which is a typical interval for Gossip [20] multi-casting protocols.
To learn the optimal policy $\pi^{*}$, the agents evaluate taking action $a$ in state $s$ following policy $\pi$ using the Q-function $Q_{\pi}(s,a)=\mathbb{E}\left[G\mid s,a,\pi\right]$. Therefore, the optimal Q-function $Q^{*}(s)=\max_{\pi}Q_{\pi}(s)$ cannot be less than the Q-function of any other policy, i.e., $Q_{\pi^{*}}(s)\geq Q_{\pi}(s)$. Deep Q-Learning (DQL) approximates the Q-function using a mapping $\phi$ of state-action pairs, allowing it to work in high-dimensional and continuous state spaces. This mapping is weighted by $\theta$ to calculate the Q-function $Q(s,a)=\theta\phi(s,a)$, which is updated according to Bellman equation: $Q(s,a)\leftarrow Q(s,a)+\alpha r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-\enspace Q(s,a)$.
The best action in a given state is the one that maximizes the Q-function for that state, i.e., $\pi(a\mid s)=\operatorname*{arg\,max}_{a}Q(s,a)$. Once the algorithm converges to $Q^{*}$, the optimal action is predicted using the optimal policy $\pi^{*}(s)=\operatorname*{arg\,max}_{a}Q^{*}(s,a)~{}\forall s\in S$. To avoid instability and divergence during training, a random set of interactions is batched from a Replay Buffer to train the Q-function network [21]. In addition, an off-policy Double-DQL [22] balances exploration and exploitation by learning an optimal policy while using an exploratory policy for action selection. It uses two networks for the Q-function, one updates every training step to evaluate the actions, and the other updates periodically for stable action selection.
After learning the optimal policy, we allow the agents to stop training, and use the learned models for inference only. This mitigates the need for unnecessary computations and leverages the simple arithmetic and array operations of lightweight inference models. The agents can then detect changes in the environment by measuring degradation in performance, which is used to trigger the fine-tuning process for their models to adapt to the new environment. Fine-tuning the learned models in a lifelong learning process can speed up convergence, improve data efficiency, and provide better generalization than learning from scratch [23]. In this work, for instance, the agents experience three significant surges in workload generation rates, causing three inference/fine-tuning epochs. Fine-tuning can also be used to initialize new agents by leveraging the expertise of existing trained agents in the network (see Fig. 5).
Figure 5: Lifelong learning for RL agents.
Exploiting previously learned knowledge $\mathcal{I}_{\mathcal{S}}$ from source task $\mathcal{S}$ to improve the learning on a related target task $\mathcal{T}$ requires the tasks to have a certain level of similarity [24]. Hence, the learned policy on the target environment $E_{\mathcal{T}}$ depends on the information from both source and target tasks, i.e., $\pi_{\theta}=\phi(\mathcal{I}_{\mathcal{S}}{\sim}E_{\mathcal{S}},\mathcal{I}_{\mathcal{T}}{\sim}E_{\mathcal{T}}):S^{\mathcal{T}}\rightarrow A^{\mathcal{T}}$. This information includes the parameters of the Q network ( $\theta_{\mathcal{T}}\leftarrow\theta_{\mathcal{S}}$) and the $x$ recent interactions of the replay buffer ($\mathcal{B}_{\mathcal{T}}\leftarrow\text{RecentExperiences}(\mathcal{B}_{\mathcal{S}},x)$).
V Performance Evaluation
This paper mainly focuses on comparing the learning of independent decision-making policies with the learning of a common policy. A common policy exists in both single-agent and MARL CTDE solutions. Hence, we evaluate our proposed fully distributed MARL approach against a common agent policy as in [2]. This highlights the advantages of training independent agents compared to a single LB agent, including autonomy, adaptability, and scalability. For instance, $AP_{2}$ in Fig. 6 can quickly learn, after a few interactions, that $F_{1}$ is not feasible to process any of its generated workloads, given its distance and limited resources. Similarly, $AP_{1}$ can quickly learn that it can use $F_{2}$ to process its compute-intensive workloads, despite its distance.
Figure 6: Benefit of independent distributed policies.
Using a common policy to learn these infeasible decisions requires more interactions than independent policies, especially in large and complex environments. A common policy explores a larger number of state-action combinations, while independent policies provide more targeted exploration for each agent by focusing on the infeasible decisions for every agent separately. Independent policies mitigate the burden of exploring vast state and action spaces, allowing for faster learning even with larger spaces. Hence, the results of this section will help show the ability of our fully independent agents to learn better distribution strategies compared to a common centralized policy.
We leverage the system parameters based on our previous contributions [18, 17, 2], which ensures consistency and relevance in our evaluation, and provides valuable insights into the effectiveness and applicability of the fully distributed MARL approach in real-world scenarios. We used TF-Agents [25] to implement the DDQL algorithm with the hyper-parameters in Table III. A fully independent agent is deployed in each AP with its policy model, Q-Network, and experience replay buffer without sharing learning information between the agents. The simulations run on a DELL G5 laptop with 16GB RAM and Intel Core i7-9750H (6 physical cores, 12MB Cache, running up to 4.5 GHz). The agents are trained using NVIDIA GeForce RTX 2070 GPU with Max-Q Design (8GB RAM). To simulate realistic Fog environments we used a Discrete-Event Simulator called YAFS [26].
TABLE III: DDQL Hyper-parameters
Parameter Values
Discount Factor $\gamma$ 0.99
Decayed Epsilon-Greedy The first 75% Training steps
Decayed Epsilon values Linearly from 100% to 1%
Replay Buffer Type Uniform
Maximum Buffer Capacity 1 Million trajectories of experience
Initial Buffer Population 10% of its maximum capacity
Buffer mini-Batch Size 50 samples of 2 steps each
Network Train Period 4 Decision steps
Target Update Period 2000 Decision Steps
Network Layers Fully connected [256, 128, 64]
Network Optimizer Adam with $2.5e^{-4}$ learning rate
Network Loss Function Huber Loss
Flat interconnected Fog topologies are generated using the Networkx library [27] as random undirected graphs resembling the Internet Autonomous System network. Nodes in the graph are then identified as Cloud, Fog nodes, or APs based on their betweenness centrality [28]. For instance, Fig. 7 shows a sample topology with a Cloud, 11 Fog nodes, and 21 APs. Fog nodes are then assigned heterogeneous compute resources, in Instructions Per Second (IPS), from evenly spaced values in the range $[10^{3},10^{5}]$. This is done inversely proportional to the number of directly connected APs to resemble unbalanced load distribution.
Figure 7: Unbalanced flat Fog topology with interconnected Fog nodes.
The Cloud, with $10^{6}$ computing resources, is connected to two Fog nodes with the highest betweenness. Fog nodes are assumed closer to IoT devices with low link latency compared to the Cloud, which incurs higher delays due to longer transmission distances and additional processing layers. To simulate the physical distances of IoT-Fog, Fog-Fog, and Fog-Cloud links, the propagation delay of those links, in seconds, is drawn from a uniform distribution in the ranges $[1,2)$, $[2,4)$, and $[10,20)$, respectively [7]. While their bandwidth, in Mbps, is drawn from a uniform distribution in the ranges $[10^{2},10^{3})$, $[10^{3},10^{4})$, and $[10^{3},10^{4})$, respectively [7]. Heavy, moderate, and light workloads require $10^{4}$, $10^{3}$, and $10^{2}$ compute instructions, respectively. The data sizes of these workload categories are $10^{3}$, $10^{2}$, and $10^{1}$ Bytes, respectively. These values were chosen relative to Fog and link resources to create a system with scarce resources relative to the demand.
Workloads are generated as a Poisson Point Process using exponential distributions to effectively model the stochastic nature of job arrivals in Fog computing environments [29]. We used three scales for the exponential distributions, i.e., $\beta=$ 100, 150, and 200, to evaluate the performance under various load intensities, i.e., high, medium, and low generation rates/frequencies, respectively. Without efficient load distribution, the high rate can easily create a system bottleneck with heavy workloads and the unbalanced structure of the environment. In episodes of 10K simulation steps, training starts with the low rate for 30K training steps. A training step spans four decision steps, each spanning a variable number of simulation steps, depending on the workload generation rate. After 30K training steps, the rate instantly increases to medium, and then to high, requiring the agents to adapt to this rapid change using transfer learning.
To evaluate the agents’ ability to reduce the number of queued jobs in each Fog node, we use the average system waiting delay as the main performance metric. We then evaluate balance and fairness by examining the similarity in waiting times and resource utilization across Fog nodes. Finally, we evaluate the system average for end-to-end workload execution delay, which includes network latency besides Fog waiting and service times. To underscore the benefits of learning independent policies over a common decision-making policy, we compare the proposed fully Distributed RL (DRL) approach against a Centralized single-agent RL (CRL) solution. We also include four baseline methods for completeness to show the significant difference between simple and intelligent load distribution solutions. These baselines include the selection of a random Fog node for each generated workload, and a Distributed Round-Robin (DRR) mechanism, where each AP assigns each workload among Fog nodes in a circular order. We also include the selection of the nearest Fog node to each workload’s AP. The final baseline is selecting the fastest Fog node for each workload based on compute requirements relative to node and link resources.
V-A Performance Analysis
The results presented in this section are averaged over 10 experiment trials with different random seeds to showcase the confidence intervals across different Fog topologies generated with the random graph generator. Using confidence intervals provides a measure of the variability in the results to assess the precision and reliability of the average values obtained from multiple experiment trials, enhancing the robustness of the findings. Due to the significant performance differences between the proposed DRL approach and the simple baseline methods, we used a logarithmic scale to show the values of DRL in the figures that compare it with baselines. The results are gathered using the highest generation rate, After training on the three rates each with 30K training steps, as it is the most difficult for the agent to learn [2]. We use an evaluation episode of 100K simulation steps, i.e., 10×10$\times$10 × longer than training episodes to offer a more reliable measure of RL effectiveness and generalization capability in real-world scenarios.
Starting with the average system waiting delay, i.e., the average waiting time in the queues of all Fog nodes, Fig. 8LABEL:sub@fig:avgWaiting1 illustrates the performance of RL approaches compared to the baseline methods. In our evaluations, CRL performs better in reducing the system waiting delay than traditional baseline methods. This highlights the effectiveness of leveraging RL-based interactive learning to optimize load distribution in Fog environments. Furthermore, transitioning to DRL yields notable performance improvements beyond the CRL. By distributing decision-making across APs, the fully distributed independent agents capitalize on localized information and dynamic adjustments to the actions of other agents, resulting in even lower system waiting delays.
(a) DRL compared to CRL and baseline methods
(b) Realistic vs. unrealistic observations
Figure 8: The average system waiting delay.
Fig. 8LABEL:sub@fig:avgWaiting2 focuses on the performance of both centralized and distributed RL solutions with respect to the frequency of observing the environment. It shows the performance of realistic observations using a Gossip protocol with 3 seconds interval between two consecutive actions. This is compared against the unrealistic assumption, commonly assumed in the literature, of readily available observations for every action. The figure highlights that realistic observations deliver slightly inferior performance compared to the optimal scenario where observations are unrealistically provided in real-time. This underscores the trade-off between realism and performance in RL-based load distribution solutions, where performance is directly related to the observation frequency. Despite the marginal performance degradation, realistic observation intervals align with practical constraints often encountered in real-world environments.
After evaluating the average waiting delay in the system, we show the variations in waiting delay among all Fog nodes, where lower values represent similar waiting times across the Fog nodes. Fig. 9LABEL:sub@fig:stdWaiting1 shows that RL-based solutions exhibit fewer variations than the baselines, with significantly better performance for DRL. By distributing RL decision-making, the agent in each AP manages the distribution of its workloads independently from other agents. This leads to a more effective load distribution considering more localized resources and adapting to the decisions of the other agents in the system. In contrast, CRL may exhibit higher variations in waiting time compared to DRL due to the centralized decision-making entity. The centralized agent struggles to efficiently adapt its decisions to the diverse workload demands from multiple APs across the heterogeneous Fog nodes.
(a) DRL compared to CRL and baseline methods
(b) Realistic vs. unrealistic observations
Figure 9: The variations in the waiting delay between Fog nodes.
Comparing DRL and CRL with respect to the frequency of observing the environment, Fig. 9LABEL:sub@fig:stdWaiting2 shows a slightly different trend when it comes to the variations in waiting delay between Fog nodes. DRL with real-time observations is the best to equalize the waiting time across all Fog nodes. Using interval-based observations, DRL still performs better than CRL but with more variability across the different trials. CRL with real-time observations, however, shows the largest variability across trials, indicating less confidence,i.e., uncertainty, on the behavior of the agents during deployment. Interestingly, CRL with interval-based observations, with slightly less equalized waiting delay across Fog nodes, has less variability across trials than real-time observations.
Fig. 10LABEL:sub@fig:stdUtilization1 shows DRL’s ability to provide fair resource utilization between Fog nodes by reducing the variation in Fog resource utilization across the network. It shows significant improvement in DRL resource utilization fairness compared to CRL and other baselines. This mitigates the probability of overloading or underloading individual Fog nodes in the system. Maximizing the utilization of available resources enhances system performance by encouraging all Fog nodes to contribute equally to the overall workload processing, relative to their capacity constraints. On the other hand, mitigating overloading enhances system stability by ensuring the maximum availability of spare resources across the network, which minimizes the probability of dropping workloads in overloaded Fog nodes in case of a sudden surge in workload generation.
(a) DRL compared to CRL and baseline methods
(b) Realistic vs. unrealistic observations
Figure 10: The variations in the module utilization between Fog nodes.
Fig. 10LABEL:sub@fig:stdUtilization2 shows a small degradation in DRL fairness after using realistic interval-based observations. On the other hand, the choice between real-time and interval-based observations does not seem to have a significant difference in CRL fairness. Unlike Fig. 9LABEL:sub@fig:stdWaiting2, Fig. 10LABEL:sub@fig:stdUtilization2 shows small variability across the different trials, indicating less confidence on agents behavior. It is important to analyze agents’ ability to provide fairness in resource utilization, similar waiting delay between all Fog nodes, and minimal average waiting time. However, we still need to observe the average end-to-end execution delay, including network links latency, Fog queue waiting time, and Fog service time for all generated workloads.
Fig. 11LABEL:sub@fig:execution1 shows the ability of RL-based solutions to significantly minimize, compared to baseline methods, the average end-to-end execution delay for all generated workloads in the system. This performance improvement is directly attributed to agents’ ability to reduce the accumulation of workloads in the queues of Fog nodes, emphasizing the critical role of reducing the waiting delay in enhancing overall system performance. As shown earlier, there is a significant difference between DRL and CRL’s ability to minimize the variations in waiting delay and resource utilization between Fog nodes. However, the total execution delay achieved by both methods is quite similar, which is again attributed to their ability to reduce the waiting delay to a similar level relative to the performance of baseline methods.
(a) DRL compared to CRL and baseline methods
(b) Realistic vs. unrealistic observations
Figure 11: The average total execution delay.
The relatively small difference in total execution delay between DRL and CRL (see Fig. 11LABEL:sub@fig:execution2) underscores the efficacy of both approaches in optimizing system performance through different load distribution strategies. However, performance difference simply amplifies with less efficient solutions over longer simulations due to the accumulation of workloads over time, leading to increased total execution delays with DRL compared to CRL. Fig. 11LABEL:sub@fig:execution2 also shows a small degradation in DRL performance using interval-based observations compared to CRL, but still performance better than CRL. The performance of CRL, on the other hand, is almost identical between real-time and interval-based observations.
In summary, the results presented in this section offer insights into the performance of our proposed fully distributed MARL solution, i.e., DRL, against the performance of a single-agent RL solution and some other baselines. The results demonstrate the superior performance of RL approaches over baseline methods, with DRL showcasing the most notable improvements, highlighting the effectiveness of distributed decision-making in optimizing workload distribution. These improvements will amplify over longer simulations due to the accumulation of workloads over time, making such intelligent solutions the best option for real-world implementations. However, it’s crucial to note that the realistic performance of such solutions highly depends on observation frequency through multi-casting protocols, such as Gossip.
VI Conclusion
This paper presents a MARL fully distributed load-balancing solution that optimizes end-to-end execution time by minimizing Fog nodes’ waiting delay. The agents in our design work independently, and can dynamically adapt to environment changes using transfer learning in a life-long learning process. Our findings demonstrate the effectiveness of the proposed distributed solution in minimizing waiting time while enhancing resource utilization fairness across heterogeneous fog resources. By leveraging distributed decision-making, MARL agents outperform centralized solutions and traditional baselines, leading to significant improvements in overall system performance. Furthermore, we evaluate a more realistic approach to observing the state of the environment through an interval-based multi-casting protocol. Considering the real cost of observing the environment has never been considered before, and instead, observations are assumed to be unrealistically available in real-time for every decision to make.
Unlike centralized load-balancing solutions, using fully distributed agents allows for a global-scale solution where groups of agents can work independently in small collaboration regions. For this reason, a primary focus for future research is to provide a real-world implementation to demonstrate the scalability and effectiveness of the proposed distributed load-balancing solution at a global scale. This can involve deploying independent agents in diverse Fog environments that span across distant geographical regions and heterogeneous network infrastructures. This research direction will bridge the gap between theoretical research and practical implementation, demonstrating the practical feasibility and effectiveness of the proposed solution at a global scale. In addition, testing with a real-world implementation can uncover potential limitations and constraints that can guide future research efforts in enhancing Fog load-balancing solutions.

### v2

@@
 Real-time Internet of Things (IoT) applications require real-time support to handle the ever-growing demand for computing resources to process IoT workloads.
-Fog Computing provides high availability of such resources in a distributed manner.
-However, these resources must be efficiently managed to distribute unpredictable traffic demands among heterogeneous Fog resources.
-This paper proposes a fully distributed load-balancing solution with Multi-Agent Reinforcement Learning (MARL) that intelligently distributes IoT workloads to optimize the waiting time while providing fair resource utilization in the Fog network.
-These agents use transfer learning for life-long self-adaptation to dynamic changes in the environment.
-By leveraging distributed decision-making, MARL agents effectively minimize the waiting time compared to a single centralized agent solution and other baselines, enhancing end-to-end execution delay.
-Besides performance gain, a fully distributed solution allows for a global-scale implementation where agents can work independently in small collaboration regions, leveraging nearby local resources.
-Furthermore, we analyze the impact of a realistic frequency to observe the state of the environment, unlike the unrealistic common assumption in the literature of having observations readily available in real-time for every required action.
-The findings highlight the trade-off between realism and performance using an interval-based Gossip-based multi-casting protocol against assuming real-time observation availability for every generated workload.
+Fog Computing provides high availability of such resources closer to IoT devices in a distributed manner.
+However, these resources must be efficiently managed to distribute unpredictable IoT traffic demands among heterogeneous Fog resources.
+To fully leverage these distributed Fog resources, their management must be done in a fully distributed manner, i.e., without any source of centralization management.
+Hence, this paper proposes a fully distributed load-balancing solution with Multi-Agent Reinforcement Learning (MARL) that intelligently distributes IoT workloads to minimize waiting delay while providing fair resource utilization in the Fog network, enhancing end-to-end execution delay.
+By leveraging distributed decision-making without centralized or inter-agent coordination, MARL agents can provide faster and better load distribution decisions with minimal overhead compared to a single centralized agent solution.
+Besides performance gain, a fully distributed solution allows for scalable Fog load balancing, where agents manage subsets of a global Fog network in smaller collaboration regions.
+To ensure global optimality, the agents learn to optimize a common objective, i.e., aiming to reduce workload accumulation in every Fog node.
+Reducing state and action spaces for each agent to a few Fog nodes in proximity makes learning easier and convergence faster.
+Additionally, learning independent policies by avoiding coordination encourages the agents to leverage Fog resources in proximity before seeking remote resources.
+In this paper, we also evaluate the impact of a realistic frequency for observing the environment instead of the common assumption of having observations readily available for every action, highlighting the trade-off between realism and performance and the solution’s applicability to real-world deployment constraints.
 Index Terms:
-Internet of Things, Fog Computing, Load Balancing, Reinforcement Learning, Multi-Agent Learning
+Internet of Things, Fog Computing, Load Balancing, Reinforcement Learning, Multi-Agent Learning.
 I Introduction
@@
 Load balancing in Fog computing environments plays a crucial role in ensuring optimal resource utilization and timely processing of IoT data; this allows for minimizing task completion time and improving overall network performance.
-Traditional load-balancing techniques often struggle with the dynamic nature of Fog environments, characterized by heterogeneous resources, limited communication bandwidth, and fluctuating workloads.
-This is where Reinforcement Learning (RL) comes into play, as it allows agents to learn optimal decision-making policies through interaction with the environment.
-However, managing Fog and IoT resources must be done at the network edge, in a distributed manner.
-This will avoid the need for distant decisions, taking full advantage of the distributed computing capabilities of Fog systems [1].
-Unlike centralized RL-based load-balancing solutions, fully distributed Multi-Agent RL (MARL) overcomes scalability limitations by allowing autonomous agents to learn from their local observations and interactions.
-These agents can be deployed in network Access Points (APs), i.e., Base Stations, to distribute the load their associated end devices generate.
-This paper presents a novel fully distributed approach for Fog load balancing using MARL.
-Our approach introduces distributed self-learning agents capable of efficiently distributing IoT load across the Fog environment.
-These agents incorporate self-adaptation mechanisms, leveraging practical lifelong transfer learning to adapt to dynamic environmental changes seamlessly.
-This distributed solution outperforms the centralized lifelong learning agent proposed in our previous work [2].
-Available MARL solutions in the literature provide collaboration between RL agents through one or more centralized components, like a centralized policy and a centralized experience replay buffer, which is not practical for a global-scale Fog load-balancing solution.
-Without fully decentralizing decision-making, the communication overhead grows exponentially with the number of agents and Fog nodes in the system, especially when the agents need to communicate with each other.
-In addition, proposing fully independent agents lowers the state and action spaces of individual agents to reduce the computational complexity of the solution.
-This is done by considering smaller sets of Fog nodes, e.g., within a certain proximity, as possible candidates for each group of agents (see Fig.
-1).
+Traditional load-balancing techniques often struggle with the dynamic nature of Fog environments, characterized by limited heterogeneous resources and fluctuating workload demands.
+Reinforcement Learning (RL) can be used to learn optimal load distribution policies by interacting with such dynamic environments.
+To take full advantage of the distributed computing capabilities of Fog systems, load distribution decisions must be provided in a distributed manner closer to IoT devices at the network edge [1].
+Fully distributed Multi-Agent RL (MARL) overcomes the scalability limitations of centralized solutions by allowing autonomous agents to independently manage smaller sets of Fog nodes in proximity instead of managing a larger network by a single agent.
+These gents incorporate self-adaptation by leveraging the practical lifelong transfer learning framework proposed in our previous work [2], allowing them to adapt to dynamic environmental changes seamlessly.
+Each agent is deployed in an IoT Access Point (AP), i.e., Base Station, to learn its own load distribution strategy for its associated IoT devices considering only the candidate Fog nodes in its region.
+Managing the same set of candidate Fog nodes by multiple agents forms a collaboration region (see Fig.
+1), where the goal is to optimize these shared resources independently by each agent.
+These collaboration regions can be isolated, but they can also be overlapped by sharing candidate Fog nodes between collaboration regions as in Fig.
+1.
 Figure 1: A simplified Fog network with overlapping collaboration regions.
-Fig.
-1 shows the formation of smaller collaboration regions, where the group of agents accessing the same set of candidate Fog nodes can form a collaboration region.
-Each region contains one or more agents that need to distribute the load between one or more candidate Fog nodes.
-In this setup, an agent can only be part of a single collaboration region.
-Candidate Fog nodes can be shared between two or more overlapping regions, while an isolated collaboration region does not share its set of candidate Fog nodes with other regions.
-Allowing the agents to independently, i.e., blindly, learn to collaborate in each collaboration region leads to the global optimization of the whole system.
-The independent agents must learn to optimize the same objective of minimizing the overall system waiting delay.
-To avoid unnecessary communication overhead, each agent observes the same state from the environment while independently taking actions and receiving rewards for those actions.
-Hence, our independent distributed agents can work without sharing their knowledge, experience, or actions, avoiding the need for a centralized coordination entity and inter-agent communication.
-In addition, a realistic multi-cast approach for observing the environment is proposed in this work, compared to unrealistic real-time observations often considered in the literature.
-This paper proposes the first fully distributed, efficient, and self-adapting MARL load distribution solution for realistic global-scale Fog Computing implementations without centralized components or inter-agent communication.
-Through empirical evaluation, we show the effectiveness of our novel truly distributed MARL-based load-balancing solution and compare it with baseline solutions and a centralized state-of-the-art RL solution.
-We also show the performance of observing the environment using realistic multi-cast intervals compared to unrealistic real-time observations.
+Considering small collaboration regions lowers the state and action spaces of individual agents, reducing the computational complexity of the problem and making the learning task easier for each agent.
+To simplify the problem further, we allow the agents to learn independent policies without collaboration, where they learn to take independent actions that fit their unique load requirements and the Fog resources closer to each agent.
+Learning separate load distribution strategies can be faster than learning a shared strategy for all agents, but requires a way to ensure global optimality while avoiding making the environment non-stationary for each agent.
+To do this in both isolated and overlapped regions, all agents learn to optimize a common objective through a common reward function, designed to reduce system-wide workload accumulation by learning to select less loaded Fog nodes.
+To avoid learning a common load distribution strategy, agents must learn in a fully distributed manner without sharing any information with other agents in the system, either through a centralized coordination entity or inter-agent communication.
+This is because coordination eventually leads to learning a common policy, either by explicitly learning a shared policy with distributed execution of that policy or by implicitly learning a common policy through parameter and experience sharing.
+Parameter sharing is done by exchanging the learned parameters of the neural network (knowledge) between agents.
+Experience sharing can be done by exchanging the actions or the full interactions in the replay buffer.
+Besides leading to a common policy, coordination causes significant communication overhead, especially with inter-agent communication with high information-sharing frequency.
+Most MARL solutions in the literature consider centralized collaboration through one or more centralized coordinators, reproducing the scalability problem we are trying to solve in the first place.
+Moreover, the overhead of observing Fog information from across the network using multiple agents is never considered in the literature.
+Researchers commonly assume this information is readily available in real-time for every agent’s action, which is impractical in network environments.
+Although this unrealistic assumption leads to theoretically optimal results, we evaluate in this work a more realistic way to observe Fog information using interval-based observations.
+We consider observation intervals of existing multi-casting protocols to evaluate a more realistic, although sub-optimal, performance of our solution.
+This paper presents a novel fully distributed, efficient, and self-adapting MARL load distribution solution for realistic global-scale Fog Computing implementations without centralized or inter-agent coordination.
+Through empirical evaluation, we show the effectiveness of our distributed MARL-based solution compared to the centralized RL solution proposed in our previous work [2].
+We also show the performance of observing the environment using realistic multi-cast intervals of 3 seconds compared to using unrealistic real-time observations.
 Hence, the key contributions of this paper can be summarized as follows:
-  • Propose fully distributed self-learning LB RL agents that efficiently distribute IoT load into a Fog Computing environment.
-  • Augment the independent distributed agents with lifelong self-adaptation to dynamic environment changes using transfer learning.
-  • Provide a global-scale solution by allowing the agents to work independently in small collaboration regions, unlike centralized architectures.
-  • Evaluate realistic interval-based collection of observations using Gossip multi-casting protocol compared to the common practice in the literature of unrealistically providing this information in real-time.
+  • Propose a scalable Fog load balancing solution for a global-scale implementation using fully distributed RL agents that blindly collaborate without coordination by learning to optimize a common objective through independent actions.
+  • Problem complexity is reduced by grouping the agents into smaller collaboration regions, with isolated or overlapped Fog nodes between the regions, to observe and manage smaller subsets of Fog nodes from the global network.
+  • These distributed agents reduce training time by incorporating a lifelong self-adaptation framework using transfer learning, allowing the agents to efficiently distribute dynamically changing IoT load across the Fog environment.
+  • Evaluate the realism-performance trade-off based on the frequency of observing Fog information from across the network, where we compare the common practice in the literature of unrealistically assuming Fog information is readily available in real-time for every action against the more realistic approach of collecting Fog information using existing interval-based multi-casting protocol, like Gossip.
 The rest of the paper is organized as follows.
@@
 [3] proposed training an agent for each Fog node to mitigate the delay of decision-making in centralized solutions.
-However, they trained all the agents offline on a centralized cloud data center to later distribute the trained models to each Fog node to make offloading decisions locally with minimal delay.
-However, training these agents using this approach, and retraining them in case of changes in the environment, requires the collection of local experience trajectories of each agent, i.e., observations, actions, rewards, and next observations.
+However, they trained all the agents offline on a centralized cloud data center to later distribute the trained models to each Fog node, making offloading decisions locally with minimal delay.
+Centralized training provides the same decision-making policy for all agents despite their varying requirements, conditions, and nearby resource availability.
+Also, centralized training, and retraining in case of changes in the environment, require collecting local experience trajectories of each agent, i.e., observations, actions, rewards, and next observations.
 Transferring such information to a centralized entity to retrain these agents is inefficient compared to local retraining, especially for highly dynamic environments that keep changing over time.
-Determining the optimal timing for retraining these agents in the centralized entity presents a critical trade-off between decision accuracy and the significant retraining overhead.
+Determining the optimal timing for retraining the agents in a centralized entity presents a critical trade-off between decision accuracy and the significant retraining overhead.
 In addition, an RL decision can only be made when the observation of the current step is available, along with the reward of the previous step.
 Observing the state, for each decision, of every Fog node in large-scale Fog networks will cause a significant delay, especially with frequently generated workloads [4].
-This delay is often not considered in the literature when proposing load-balancing solutions with real-time decision-making.
-Hence, the effectiveness of these solutions in practice can not be guaranteed without considering the time needed to transfer Fog load provisioning information to all RL agents.
+This delay is exponentially proportional to the number of agents in a distributed solution, making it impractical in real-world network deployment either with a centralized or a distributed solution.
+This delay is often not considered in the literature for Fog load-balancing solutions, assuming observations are readily available for every action at no cost.
+Without considering the time needed to transfer Fog load provisioning information to all RL agents, the effectiveness of these solutions in practice can not be guaranteed.
 Gao et al.
@@
 In their design, each edge server is regarded as an actor agent while still considering a centralized critic, where the critic is a neural network used to evaluate the actions of actor networks [6].
-This centralized critic facilitates the coordination of offloading policies among those distributed actors.
-Besides not being fully distributed, the authors assumed that observations are readily available when the agents need to make assignment decisions.
-This is a common unrealistic assumption in the literature, where the delay of exchanging Fog status information to every agent is often ignored.
+This centralized critic facilitates the coordination of offloading policies among those distributed actors, leading to a common policy.
+Besides not being fully distributed, the authors assumed that observations are readily available when the agents need to make assignment decisions, where the delay of exchanging Fog status information to every agent is ignored.
 Baek et al.
@@
 Those distributed agents are located in APs and they cooperatively learn to solve the joint link and server load-balancing problem.
-The policy of the distributed cooperative agents is optimized using parameter-sharing and distributed execution; the agents communicate among themselves while evaluating a common value function.
+A common policy for the distributed cooperative agents is optimized using parameter-sharing and distributed execution; the agents communicate among themselves while evaluating a common value function.
 The authors assume that updated policy parameters are shared between the agents by exchanging communication messages over limited bandwidth channels.
@@
 [8] proposed a MARL framework to improve the execution delay of heterogeneous Fog services using multiple actors and a centralized critic.
-They deployed distributed agents across different Fog nodes to work in parallel, i.e., brokers that make service offloading decisions.
-Then, they train a centralized service offloading model, i.e., the learner of the target policy, using the experience trajectories generated by the distributed actors.
-But again, they did not consider the delay caused by exchanging such information in a large-scale environment with high request rates.
-In addition, they did not consider the delay when updating the common value function of each actor from the centralized critic network.
+They deployed distributed agents (they call them brokers) that work in parallel to make service offloading decisions across different Fog nodes.
+A centralized service offloading model, i.e., the learner of the target policy, is trained using experience trajectories generated by the distributed actors.
+But again, they did not consider the delay caused by exchanging such information in a large-scale environment, especially when request rates are high.
+In addition, they did not consider the delay when updating the value function of each actor from the common centralized critic network.
 Wei et al.
@@
 Like most MARL-based solutions in the literature, they used the approach of centralized training with decentralized execution (CTDE).
-They used the attention mechanism to share the agents’ parameters with a joint critic network.
-This joint network incorporates the learning process of distributed gated actors that exploit the local computational potentials.
+They share the agents’ parameters with a joint critic network using the Attention machine learning mechanism, leading to a joint network that incorporates the learning process of distributed gated actors that exploit local computational potentials.
+Likewise, Suzuki et al.
+[10] used CTDE to send learning samples from each agent to a centralized training entity, including task information, network link utilization, and Fog resources utilization.
+They unrealistically assumed all this information readily available in real-time for all agents when making offloading decisions.
 Zhang et al.
@@
 During distributed execution, on the other hand, only local environment information is needed for each edge node to make the request scheduling decision.
-However, the authors did not discuss the scope of the local information, and the delay associated with transmitting such information over the network.
-In a large-scale network, this delay is not tolerable for real-time decisions required by real-time applications.
+However, the authors did not discuss the scope of local information and the delay associated with transmitting such information over the network.
+In a large-scale network, such a delay is not tolerable for real-time decisions required by real-time applications.
 Lu et al.
@@
 However, offline execution does not allow the agents to dynamically adapt to environment changes.
-In addition, their solution will not scale to larger networks as all their agents must observe the global state of the environment, which are assumed, unrealistically, readily available for every decision to be made.
+In addition, their solution will not scale to larger networks as all their agents must observe the global state of the environment, which is assumed, unrealistically, readily available for every decision to be made.
 CTDE was also used by Gao et al.
@@
 Then, the agents independently make their decisions using their policies distributively.
-Likewise, Suzuki et al.
-[14] used CTDE to send learning samples from each agent to a centralized training entity to avoid the non-stationary problem of independently trained agents.
-A non-stationary environment is when the new state depends on the joint action of multiple agents instead of a single agent’s behavior.
-To learn and make offloading decisions, the agents in these approaches observe task information, network link utilization, and Fog resources utilization, which are all unrealistically assumed available in real-time for all agents decisions.
 Mangiaracina et al.
@@
 They estimate the global information in a centralized model to reduce the communication overhead caused by broadcasting this information from each agent to the centralized critic network.
-They showed a 20% gap in performance between their approach and the optimal solution but with only half the decision-making time, where the optimal solution is a centralized decision-making method having the states and actions available for all agents.
-Most existing MARL solutions adopt a centralized training approach, no fine-tuning for dynamic adaptation, and unrealistic assumptions for observation availability (see Table I).
-The centralized component in these solutions raises several challenges, especially in large-scale environments, including computational complexity, exploration of larger state and action spaces, and the communication overhead of inter-agent exchanges.
-Despite the efforts to mitigate these challenges in CTDE-based solutions, a truly distributed solution remains preferable to enhance scalability and decision-making time.
-With true distribution of decision-making, each agent makes decisions independently, without having access to the state of other collaborating agents in the system.
+They showed a 20% gap in performance between their approach and the optimal solution but with only half the decision-making time, where the optimal solution is a centralized decision-making method having the actions of each agent and the actual state information of each Fog node (not estimated) available for all agents unrealistically in real-time.
+Based on the literature discussed above, most existing MARL solutions adopt a centralized training approach and rely on unrealistic assumptions regarding collecting observations (see Table I).
+The centralized component in these solutions raises several challenges, especially in large-scale environments, including computational complexity, exploration of larger state and action spaces, and the communication overhead due to exchanging training information with the centralized entity.
+A truly distributed solution remains preferable to enhance scalability and decision-making time, where agents can learn and make decisions independently without centralized or inter-agent coordination.
+Allowing the agents to learn independent policies based on the unique requirements of every agent can provide faster convergence compared to learning a common policy for all the agents.
 TABLE I: Comparison of state-of-the-art MARL-based Fog load balancing solutions.
@@
 Our fully distributed approach Distributed Lifelong learning None Interval-based observations
-This paper proposes a fully distributed solution that efficiently coordinates decision-making among independent agents without reliance on centralized components or inter-agent communication.
-This truly distributed MARL framework effectively handles the complexities of Fog environments, scales well to a global-scale solution, and minimizes communication overhead and decision latency.
-In addition, this work presents the first attempt to evaluate a realistic way to observe the environment and the reward of the actions of each agent based on a multi-cast Gossip protocol.
+To fill this gap in the literature, this paper proposes a fully distributed solution using independent agents that efficiently distribute IoT load in Fog networks without reliance on centralized components or inter-agent communication.
+Besides mitigating communication overhead, avoiding coordination between the agents leads to learning separate decision-making policies, which is easier than learning a common load distribution strategy.
+These independent agents learn to reduce workload accumulation in Fog nodes by receiving a reward for every load assignment decision representing the accumulated load in the selected Fog node.
+Optimizing this common objective encourages each agent to leverage nearby resources before seeking remote resources, mitigating the effect of the non-stationarity problem of independently trained agents.
+This reward function allows for a global-scale implementation when forming smaller groups of agents each managing a small subset of nearby Fog nodes.
+This grouping reduces state and action spaces for the agents, making learning easier and state observation overhead smaller.
+To reduce the overhead of observing Fog information further, we present the first attempt to evaluate collecting state observations and rewards using an interval-based Gossip multi-casting protocol.
+This approach is more realistic than assuming such information is readily available immediately before every decision.
+This evaluation explores the trade-off between optimality and applicability of the solution in real-world networks, leading to a deployment-ready solution.
 III The Fog System Model
@@
 Table II shows the definitions of the mathematical symbols used throughout this paper.
-We reproduce the realistic Fog environments used in our previous works [17, 18] to compare the proposed fully distributed multi-agent solution with the centralized single-agent solution in [17].
+We reproduce the realistic Fog environments used in our previous works [17, 18] to compare the proposed fully distributed (multi-agent) solution with the centralized (single agent) solution in [17].
 TABLE II: List of mathematical notations.
@@
 $N$ Set of AP, Fog, and Cloud nodes ($N^{AP},N^{F},N^{C}$)
+$N^{F_{C}}$ Candidate Fog nodes for the agents in a given region
 $IPT_{n}$ Compute capabilities of node $n$ (Instructions/timestep)
@@
 $\mathbb{Q}_{n}$ The number of queued jobs in node $n$
-$s_{t}$ The observed state at step $t$, from a set of possible states $S$
-$a_{t}$ The agent’s action at step $t$, from a set of possible actions $A$
+$s_{t}$ State at step $t$, from a set of possible states $S$
+$a_{t}$ Action at step $t$, from a set of possible actions $A$
 $r_{t}$ The agent’s immediate reward at step $t$
@@
 Load-balancing agents are deployed in these APs, where each agent makes its own decisions to distribute the load of its associated IoT devices between the Fog nodes in its collaboration region, i.e., its candidate Fog nodes.
-Agents in each collaboration region share a set of candidate Fog nodes, requiring the agents to optimize the resources of all candidate Fog nodes collaboratively.
-To be fully distributed, the agents must collaborate blindly, i.e., without sharing their experience or decision-making policies.
-The Fog network is defined by the set of $N=n_{1},n_{2},\cdots,n_{z}$ nodes described by their compute $IPT_{n}$ and memory $RAM_{n}$ resources [17, 18].
+A group of agents in each collaboration region manages a set of candidate Fog nodes, requiring the agents to jointly optimize the resources of all candidate Fog nodes in that region.
+To be fully distributed, the agents must collaborate blindly without exchanging any kind of coordination information, like experiences, actions, or policies parameters.
+In our simulated environments, the Fog network is defined by the set of $N=n_{1},n_{2},\cdots,n_{z}$ nodes described by their compute $IPT_{n}$ and memory $RAM_{n}$ resources [17, 18].
 The Nodes, including $N^{AP}$ AP nodes, $N^{F}$ Fog nodes, and $N^{C}$ Cloud nodes, are connected through $L=l_{1},l_{2},\cdots,l_{z}$ bidirectional links.
-Each link is characterized by the pair ($n_{i}$, $n_{j}$) that it connects, its bandwidth $BW_{l}$, and its propagation delay $PR_{l}$.
-The network has a set of distributed applications $DA=da_{1},da_{2},\cdots,da_{z}$ that simultaneously run in the system; each application has a set of modules $M_{da}$ and dependency messages $D_{da}$ between these modules, i.e., workloads or requests.
-Nodes and network links have heterogeneous resources and must serve three different workload categories $|W|$ each with different compute requirements, i.e., light, medium, and heavy workloads.
-Workloads are generated as a Poisson Point Process using exponential distributions with three different scale parameters $\beta$ to simulate high, medium, and low workload generation rates, allowing for dynamic representation of job arrivals in our Fog environment.
+Each link is characterized by the pair of nodes ($n_{i}$, $n_{j}$) that it connects, its bandwidth $BW_{l}$, and its propagation delay $PR_{l}$.
+A set of distributed applications $DA=da_{1},da_{2},\cdots,da_{z}$ simultaneously run in the system; each application has a set of modules $M_{da}$ and dependency messages $D_{da}$ between these modules, i.e., workloads or requests.
+Nodes and network links have heterogeneous resources and must serve different workload categories $|W|$ each with different compute requirements, i.e., light, medium, and heavy workloads.
+Workloads are generated as a Poisson Point Process using exponential distributions with three different scale parameters $\beta$ to simulate high, medium, and low workload generation rates.
+This allows for a dynamic representation of job arrivals in our Fog environment, where small $\beta$ values increase the probability of having smaller time intervals between subsequent workloads, leading to higher generation rates.
 Nodes and links are modeled with M/M/1 queuing models to handle the generated workloads, giving the ability to provide insights into performance metrics such as queue lengths and waiting times.
-The agents must learn to minimize the number of queued jobs $\mathbb{Q}_{n}$ in the queue of each Fog node $n$ to minimize the workload waiting delay, reducing the end-to-end execution delay over the system [17].
-To optimize this objective, each agent learns a policy to select the best Fog node for each generated workload.
-To do this, each agent in a collaboration region observes the number of jobs queued in every candidate Fog node in that region.
-These observations represent the state of the environment at each step, and it is the same for all agents in a collaboration region.
-This state is used to select an action $a$ from the set of possible actions $A$, corresponding to the set of candidate Fog nodes $N^{F}$ in that region.
-Each agent learns to select the action that maximizes its performance measure, i.e., the reward, which is evaluated at the end of every decision step.
-For each agent, the reward is the number of jobs queued in the Fog node selected in its previous action.
-To reduce the average system waiting delay, this number is negated to be maximized by the RL algorithm.
-This reduces the accumulation of workloads in Fog nodes, lowering overload probability and overall execution delay.
+To minimize the waiting time before being served by Fog nodes, the agents must learn to reduce the number of queued jobs $\mathbb{Q}_{n}$ in the queue of each Fog node $n$.
+Reducing workload accumulation in each Fog node reduces Fog overload probability and the end-to-end execution delay of IoT application loops [17].
+To optimize this objective by independent agents, the agents need to learn independent policies that can balance the generated load between Fog nodes.
+To do that without exchanging coordination information, the agents observe the number of jobs queued in each Fog node in their region (as environment state $s$) to select the Fog node with the smallest load (as action $a$).
+Each agent takes its actions, from the set of possible actions $A$ corresponding to the set of candidate Fog nodes $N^{F}$ in its region, independently from all other agents.
+To reduce waiting time collaboratively by all agents, they learn to minimize the common objective of reducing workload accumulation in all Fog nodes.
+Therefore, the reward $r$ that evaluates their most recent action is defined by the load in the most recently assigned Fog node by that action, encouraging the agents to select the least loaded Fog nodes collaboratively.
+This rewarding mechanism allows the agents to jointly reduce workload accumulation in the Fog nodes of their region without sharing any collaboration information.
 IV Distributed RL Load Balancing
 A fully independent Double Deep Q-Learning (DDQL) agent is deployed in each of the $N^{AP}$ APs of the Fog network.
-The state, action, and reward function of every agent at every decision step $t$ are defined as follows:
-  • State: $s_{t}=(w_{t},\mathbb{Q}_{n,t}:\forall n\in N^{CF})$, where $w$ represents the category of the workload to be assigned and $\mathbb{Q}_{n}$ represents the number of queued jobs in candidate Fog node $n$.
-  • Action: $a_{t}\in A$ is the decision to assign the current workload to the best candidate Fog node from the set of candidates considered for that agent.
-  • Reward: $r_{t}=-\mathbb{Q}_{a_{t-1},t}$ represents the number of queued jobs in the Fog node previously selected by that agent, i.e, the reward at decision step $t$ is the performance measure of the action taken at decision step $t-1$ (it is part of the observed state, and thus, does not need to be transmitted separately to each agent).
-The negation is used to represent the reward as a cost to be minimized instead of a utility to be maximized.
+The agents in a given collaboration region are designed to manage the set of candidate Fog nodes $N^{F_{C}}$, i.e., a subset of Fog nodes from the larger network, in that region only.
+Hence, the size of state and action spaces can vary in different regions based on the number of candidate Fog nodes in each region, giving flexibility in forming those regions based on the needs of Fog service providers.
+This limits state and action spaces, and immediate rewards, to be collected from a smaller number of nearby Fog nodes, reducing the overhead of observing Fog queue information before every workload assignment decision.
+Hence, the state, action, and reward function of every agent in a given region, at every decision step $t$, are defined as follows:
+  • State: $s_{t}=(w_{t},\mathbb{Q}_{n,t}:\forall n\in N^{F_{C}})$, where $w$ is the number of compute instructions required by the workload to be assigned (representing its category) and $\mathbb{Q}_{n}$ is the number of queued jobs in each candidate Fog node $n$.
+  • Action: $a_{t}\in A$ is the decision to assign the current workload to a candidate Fog node from the set of candidate Fog nodes in the region of that agent.
+  • Reward: $r_{t}=-\mathbb{Q}_{a_{t-1},t}$ is the number of queued jobs in the Fog node previously selected by that agent, i.e, the reward at decision step $t$ is the performance measure of the action taken at decision step $t-1$.
+Because the value of the reward is part of the observed state, it does not need to be transmitted separately to each agent.
+This value is then negated to represent the reward as a cost to be minimized instead of a utility to be maximized.
 The accumulation of these rewards is the expected discounted return $G=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, where $\gamma$ defines the importance of future rewards compared to recent ones.
-In centralized single-agent solutions [2, 17], the state space consists of the number of queued jobs from all Fog nodes at each decision step, i.e., $\mathbb{Q}_{n,t}:\forall n\in N^{F}$.
-The centralized agent also requires the identity of the AP connected to the source IoT device, i.e., workload from $n^{AP}_{t}$ to be assigned at decision step $t$.
-When shifting to a distributed multi-agent solution in large-scale networks, considering all Fog nodes for every agent can unnecessarily increase the state and action spaces of each agent.
-However, it is sufficient to design the agents with feasible subsets of Fog nodes within a certain proximity.
-This leads to creating multiple groups of agents, each observing the queues of their own set of candidate Fog nodes $\mathbb{Q}_{n,t}:\forall n\in N^{CF}$.
-In addition to increasing the state and action spaces, a centralized agent causes a significant delay for each workload assignment decision due to the three-way communication between the centralized agent and the workload AP (see Fig.
+Giving importance to future rewards encourages the agents to learn a load distribution strategy that guarantees minimizing load accumulation in all Fog nodes over the long term.
+Our reward function enables the distribute agents to independently learn different load distribution strategies that keep the load in all Fog nodes minimal at all times.
+Each agent can discover its best strategy by considering its nearby Fog nodes before selecting remote Fog nodes in its region; this can be independently learned faster than learning a common policy for all agents.
+Hence, these independent agents can converge faster than single-agent solutions, as well as multi-agent solutions that lead to a common policy, i.e., explicitly and implicitly learn a common policy with CTDE and parameter/experience sharing, respectively.
+Also, learning independent policies allows us to reduce the state and action spaces by managing smaller subsets of Fog nodes, forming a collaboration region for each group of agents.
+Moreover, distributed solutions can provide near real-time decisions for every generated workload by avoiding the three-way communication overhead of centralized solutions (see Fig.
 2LABEL:sub@fig:Topology_WF_C).
-For every generated workload in centralized load distribution solution, an assignment request is sent to the centralized agent along with the workload category and the identifier of its AP.
-After receiving this request, the centralized agent predicts the best Fog node for that workload using the provided information and the observed number of queued jobs in each Fog node in the system.
-The agent then responds to the requesting AP with the identifier of the selected Fog node.
-Finally, the AP assigns and sends the workload to the selected Fog node and waits for the next generated workload.
 (a) Centralized agent
@@
 Figure 2: Workflow of workload distribution.
-With a distributed load balancing solution, however, each agent has this prediction capability locally, and hence, each agent makes its own assignment decisions (see Fig.
+In centralized solutions, a request for every generated workload is sent to a distanced agent with the category of the workload and the identifier of its AP.
+The agent notifies the AP of the suitable Fog node to serve that workload based on workload category, AP identifier, and the load in each Fog node in the network.
+Only then, the AP can send the workload to the selected Fog node.
+With distributed solutions, however, agents make their own assignment decisions locally (see Fig.
 2LABEL:sub@fig:Topology_WF_D).
-As soon as the AP receives a workload, it predicts the best candidate Fog node for that workload by providing the workload category $w$ and the number of queued jobs in each candidate Fog node.
-The workload is then immediately assigned and transmitted to the selected Fog node, minimizing the decision delay compared to a centralized solution.
-In addition to the decision-making delay induced by the agent’s location, i.e., centralized vs.
-distributed, observing the number of queued jobs in each Fog node causes a significant delay, often neglected in the literature.
-In a centralized solution, each Fog node sends its queue information to the centralized agent as in Fig.
+The workload is then immediately sent to the selected Fog node, minimizing the decision delay compared to a centralized solution.
+The time needed to observe Fog load information and the associated overhead of this process is often neglected in the literature [19].
+With a single-agent solution, all Fog nodes send their load information to that agent for every generated workload, as in Fig.
 3LABEL:sub@fig:Topology_info_C.
-In a distributed solution, however, each agent/AP needs to receive this information from all candidate Fog nodes (see Fig.
+Although our distributed agents are grouped to manage smaller subsets of Fog nodes in proximity, this overhead is still significant since every agent needs to collect this information from every candidate Fog node in their region (see Fig.
 3LABEL:sub@fig:Topology_info_D).
-To do this in a real-world environment, Fog nodes need to multi-cast their queue information, introducing a significant delay compared to a centralized solution.
+This makes MARL observation communication overhead, for every decision, exponentially proportional to the number of Fog nodes and agents in each region.
 (a) Centralized agent
@@
 Figure 3: Transmission of Fog queue information.
-Frequently multi-casting the observations, i.e., for every generated workload, in large-scale networks consumes a significant amount of network bandwidth.
-Hence, providing this information in real-time for every workload assignment decision becomes infeasible, especially with a large number of Fog nodes and IoT devices, and high workload generation rates.
-Therefore, addressing the frequency of transferring this information between the agents becomes essential.
-In addition, dividing large-scale networks into smaller subsets of collaboration regions limits the transfer of this information to fewer agents within a smaller geographical scope, thereby reducing the communication overhead.
-A distributed load balancing solution allows the network to be clustered into smaller, possibly overlapping, regions based on geographical proximity between Fog nodes and APs.
-This significantly reduces the state and action spaces by only considering feasible Fog nodes for the agents in each collaboration region, which helps reduce the computational complexity of the solution.
-The agents in each region will then collaborate within their respective boundaries to optimize the resources of their feasible candidate Fog nodes.
-Unlike centralized load balancing solutions, where all Fog nodes must be included in the observation of the centralized agent, whether they are feasible options for all APs or not.
-In this paper, we evaluate our fully distributed solution on a single collaboration region, where all agents collaborate to optimize the resources of the candidate Fog nodes in their region.
-To show the scalability of this solution to a larger network with multiple collaboration regions, these agents must be able to independently collaborate to minimize the number of queued jobs in each candidate Fog node.
-To learn independently, the agents must collaborate blindly, i.e., without knowing the policies and actions of the other agents in the system.
+Besides consuming bandwidth, observations can not be readily available before every decision due to communication delays, especially in large-scale networks with high generation rates.
+A more practical approach for deployment-ready solutions is to multi-cast such information at predefined intervals.
+This might lead to sub-optimal decisions due to using the same observation to make multiple assignment decisions, creating a non-stationary environment for the independent agents due to these outdated observations.
+This problem does not exist with centralized and coordination-based RL because the new environment state depends on a common policy behavior, not on the joint (unobserved) actions of independent agents.
+Being unrealistic to observe the real-time load on each Fog node before every decision, makes it more practical to consider a single observation for several actions within a given interval.
+Increasing observation frequency, using the smallest interval of existing multi-casting protocols (3 seconds for Gossip [20]), reduces the impact of non-stationarity.
+In addition, dividing large-scale networks into smaller subsets of collaboration regions limits the transfer of Fog load information to fewer agents within a smaller geographical scope, thereby reducing the communication overhead.
+These regions can be designed based on the requirements of Fog service providers with the possibility of creating overlapped regions by sharing Fog resources between them.
+Managing fewer Fog nodes in proximity reduces state and action spaces for each agent and avoids the need to manage infeasible remote Fog nodes in large-scale networks.
+To ensure optimal load distribution in overlapped regions, the agents must be able to blindly learn independent load distribution strategies without knowing the policies and actions of the other agents.
+By independently learning to select the least loaded Fog node, our agents minimize the load on all Fog nodes in their regions, including shared Fog nodes.
 Lemma IV.1.
-Given two non-overlapping sets of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\emptyset$ in two collaboration regions.
-Each region is managed by a group of agents that blindly collaborate to optimize the resources of their candidate Fog nodes without knowing the actions of other agents in the system (see Fig.
+Given two non-overlapping sets of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\emptyset$ in two collaboration regions (see Fig.
 4).
-If the two regions share a subset of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\mathcal{F}_{S}$, the agents of each region can still blindly collaborate to optimize the resources of their set of candidate Fog nodes and the shared set of candidate Fog nodes.
+The resources in each region are optimized by a group of agents that blindly collaborate without knowing the actions of other agents in that region.
+If both regions share a subset of candidate Fog nodes $\mathcal{F}_{1}\cap\mathcal{F}_{2}=\mathcal{F}_{S}$, the agents of both regions can still blindly collaborate to optimize local and shared Fog resources.
 Proof.
-Assuming the agents in an isolated region can efficiently manage the resources of their candidate Fog nodes by blindly collaborating without the need to observe the actions, nor the mechanism of making those actions, of other agents in their collaboration group.
-This assumption holds for any number of isolated regions each with their own set of candidate Fog nodes.
-Now, assume that a region $\mathcal{R}_{1}$ with $\mathcal{F}_{1}$ candidate Fog nodes share $\mathcal{F}_{S}$ candidate Fog nodes with region $\mathcal{R}_{2}$ (see Fig.
+Assuming the agents in an isolated region can blindly optimize their isolated Fog resources without knowing the actions of the other agents in their region, nor the mechanism of making those actions.
+Now, assume that a region with $\mathcal{F}_{1}$ candidate Fog nodes share $\mathcal{F}_{S}$ candidate Fog nodes with a region having $\mathcal{F}_{2}$ candidate Fog nodes (see Fig.
 4).
-Suppose there exists an agent in $\mathcal{R}_{2}$ that does not collaborate with the agents in its group leading to inefficient management of the Fog resources in $\mathcal{F}_{2}$, including those in $\mathcal{F}_{S}$.
+Suppose an agent in one of the regions does not collaborate with the other agents, leading to inefficient utilization of Fog resources in its region, including $\mathcal{F}_{S}$ resources.
 This implies that there is a breakdown in the collaborative optimization process.
-However, this contradicts our initial assumption that within each region, the agents efficiently collaborate to optimize the resources of their candidate Fog nodes.
-Therefore, we conclude that the agents of both regions will collaborate to optimize the shared resources in $\mathcal{F}_{S}$ as well as those of their own.
-Therefore, the lemma holds.
+However, this contradicts our initial assumption that within each region, the agents can blindly optimize the resources of their candidate Fog nodes.
+Therefore, we conclude that the agents of both regions will collaborate to optimize both local and shared resources.
+Thus, the lemma holds.
 ∎
 Figure 4: Overlapping vs. non-overlapping sets of candidate Fog nodes for two coordinating regions.
-In this work, we provide empirical evidence under a realistic setup supporting the ability of our independent distributed RL agents to optimally distribute the load within a collaboration region.
-Each agent acts according to the current environment state, and the learned policy $\pi$.
-In real-world environments, it is impractical to observe the state of every candidate Fog node for every workload assignment decision due to the network delay, which is intolerable for real-time applications.
-Observing the state of all Fog nodes in large-scale networks necessitates waiting for the state of the last node to transmit its load information to complete the environment observation.
-Hence, the effect of the observation delay is impacted by the scale of the Fog system, the number of IoT devices, and the rate of workload generation.
-In the literature, however, observations are always assumed to be immediately available for the agent at each decision step [19], rendering those solutions impractical for real-world implementations.
-To provide a practical solution with real-world constraints, we use a realistic approach to observe the number of jobs queued in each Fog node.
-In this approach, the agents collect observations using interval-based multi-casting protocols, i.e., sending queue information on predetermined intervals.
-Using the same observation for multiple decision steps will lead to sub-optimal solutions, depending on the frequency of collecting the observations.
-Hence, the smallest practical interval for the multi-casting protocol must be considered to transfer the queue information to all agents.
-In this work, for instance, Fog nodes transfer their queue information every 3 seconds, which is a typical interval for Gossip [20] multi-casting protocols.
+This paper provides empirical proof supporting the ability of our agents to blindly learn independent load distribution strategies that will jointly optimize the Fog resources in their isolated region without any coordination between these agents.
+Validating the efficiency of our fully distributed solution under a realistic setup of one isolated region proves the solution’s applicability in a global-scale multi-region network with isolated and overlapped collaboration regions.
+A fully distributed solution requires each agent to independently learn an efficient load distribution strategy, i.e., policy $\pi$, based on the category of the workload to be assigned and the current load in each Fog node in the region.
+Learning independent policies to optimize a common objective, i.e., selecting the least loaded Fog node, blindly encourages each agent to choose nearby Fog nodes before considering distanced ones.
 To learn the optimal policy $\pi^{*}$, the agents evaluate taking action $a$ in state $s$ following policy $\pi$ using the Q-function $Q_{\pi}(s,a)=\mathbb{E}\left[G\mid s,a,\pi\right]$.
@@
 It uses two networks for the Q-function, one updates every training step to evaluate the actions, and the other updates periodically for stable action selection.
-After learning the optimal policy, we allow the agents to stop training, and use the learned models for inference only.
-This mitigates the need for unnecessary computations and leverages the simple arithmetic and array operations of lightweight inference models.
-The agents can then detect changes in the environment by measuring degradation in performance, which is used to trigger the fine-tuning process for their models to adapt to the new environment.
-Fine-tuning the learned models in a lifelong learning process can speed up convergence, improve data efficiency, and provide better generalization than learning from scratch [23].
-In this work, for instance, the agents experience three significant surges in workload generation rates, causing three inference/fine-tuning epochs.
-Fine-tuning can also be used to initialize new agents by leveraging the expertise of existing trained agents in the network (see Fig.
+To mitigate the computational cost of continuous training in a practical deployment, each agent adopts our lifelong learning framework previously proposed in [2] (see Fig.
 5).
+This framework can be used to initialize new agents in a simulation, or by leveraging the expertise of existing agents in the network, before fine-tuning them in their intended deployment environment, mitigating the impact of bad actions when training from scratch in real networks.
+With this framework, the agents can stop training when they converge to the optimal policy (performance saturation) and only retrain in case of significant performance degradation due to environmental changes.
+When convergence is detected, i.e., when the average reward does not improve further, the agents extract lightweight inference models to provide inference using only simple arithmetic and array operations.
 Figure 5: Lifelong learning for RL agents.
-Exploiting previously learned knowledge $\mathcal{I}_{\mathcal{S}}$ from source task $\mathcal{S}$ to improve the learning on a related target task $\mathcal{T}$ requires the tasks to have a certain level of similarity [24].
-Hence, the learned policy on the target environment $E_{\mathcal{T}}$ depends on the information from both source and target tasks, i.e., $\pi_{\theta}=\phi(\mathcal{I}_{\mathcal{S}}{\sim}E_{\mathcal{S}},\mathcal{I}_{\mathcal{T}}{\sim}E_{\mathcal{T}}):S^{\mathcal{T}}\rightarrow A^{\mathcal{T}}$.
-This information includes the parameters of the Q network ( $\theta_{\mathcal{T}}\leftarrow\theta_{\mathcal{S}}$) and the $x$ recent interactions of the replay buffer ($\mathcal{B}_{\mathcal{T}}\leftarrow\text{RecentExperiences}(\mathcal{B}_{\mathcal{S}},x)$).
+Changes in the environment can then be detected by observing a significant degradation in the average reward, i.e., an increase in the average load accumulation in the region, triggering the need for agents to perform a corrective measure to adapt to those changes.
+The agents are then fine-tuned using transfer learning to speed up convergence, improve data efficiency, and provide better generalization than learning from scratch [23].
+Hence, policy network parameters are transferred from source to target task ($\theta_{\mathcal{T}}\leftarrow\theta_{\mathcal{S}}$) along with $x$ recent replay buffer interactions ($\mathcal{B}_{\mathcal{T}}\leftarrow\text{RecentExperiences}(\mathcal{B}_{\mathcal{S}},x)$), providing fast and efficient lifelong adaptation to changes [2].
+Since effective transfer learning requires source and target tasks to share a certain level of similarity [24], we introduce changes only through significant surges in workload generation rates.
+Using transfer learning allows the agents to learn robust policies using information $\mathcal{I}$ from both source and target environments, i.e., $\pi_{\theta}=\phi(\mathcal{I}_{\mathcal{S}}{\sim}E_{\mathcal{S}},\mathcal{I}_{\mathcal{T}}{\sim}E_{\mathcal{T}}):S^{\mathcal{T}}\rightarrow A^{\mathcal{T}}$.
 V Performance Evaluation
-This paper mainly focuses on comparing the learning of independent decision-making policies with the learning of a common policy.
-A common policy exists in both single-agent and MARL CTDE solutions.
-Hence, we evaluate our proposed fully distributed MARL approach against a common agent policy as in [2].
-This highlights the advantages of training independent agents compared to a single LB agent, including autonomy, adaptability, and scalability.
-For instance, $AP_{2}$ in Fig.
-6 can quickly learn, after a few interactions, that $F_{1}$ is not feasible to process any of its generated workloads, given its distance and limited resources.
-Similarly, $AP_{1}$ can quickly learn that it can use $F_{2}$ to process its compute-intensive workloads, despite its distance.
-Figure 6: Benefit of independent distributed policies.
-Using a common policy to learn these infeasible decisions requires more interactions than independent policies, especially in large and complex environments.
-A common policy explores a larger number of state-action combinations, while independent policies provide more targeted exploration for each agent by focusing on the infeasible decisions for every agent separately.
-Independent policies mitigate the burden of exploring vast state and action spaces, allowing for faster learning even with larger spaces.
-Hence, the results of this section will help show the ability of our fully independent agents to learn better distribution strategies compared to a common centralized policy.
-We leverage the system parameters based on our previous contributions [18, 17, 2], which ensures consistency and relevance in our evaluation, and provides valuable insights into the effectiveness and applicability of the fully distributed MARL approach in real-world scenarios.
+This paper compares learning independent decision-making policies with learning a common policy, where a common policy results from single-agent RL, CTDE MARL, or even distributed MARL training with parameter or experience sharing.
+Hence, we compare the performance of learning independent policies by our proposed fully distributed MARL approach against learning a common policy by a single agent as in [2].
+This comparison will highlight the advantages of training independent agents compared to a single LB agent, especially with convergence speed, performance, and scalability.
+To show why learning independent policies is faster than learning a common policy, Fig.
+6 shows a scenario where considering the remote rich resources of Fog node $F_{2}$ is beneficial to the agent in $AP_{1}$ while considering the remote limited resources of Fog node $F_{1}$ is bad for the agent in $AP_{2}$.
+Both agents should learn to use nearby Fog resources while carefully considering only feasible remote Fog resources.
+Using a common policy to learn a different strategy for each $AP$ is harder, more complex, and requires more training than learning an independent policy for each $AP$.
+Learning a policy for each agent independently allows every agent to find its optimal load distribution strategy faster by avoiding the complexity of learning a global policy that satisfies the needs of all $APs$.
+Figure 6: Learning a common policy vs. independently learned policies.
+The system parameters of the experiments in this work are based on the experiments in our previous contributions [18, 17, 2], ensuring consistency and relevance in our evaluation.
 We used TF-Agents [25] to implement the DDQL algorithm with the hyper-parameters in Table III.
-A fully independent agent is deployed in each AP with its policy model, Q-Network, and experience replay buffer without sharing learning information between the agents.
+A fully independent agent is deployed in each AP with its policy model, Q-Network, and experience replay buffer without sharing any information between the agents.
 The simulations run on a DELL G5 laptop with 16GB RAM and Intel Core i7-9750H (6 physical cores, 12MB Cache, running up to 4.5 GHz).
 The agents are trained using NVIDIA GeForce RTX 2070 GPU with Max-Q Design (8GB RAM).
-To simulate realistic Fog environments we used a Discrete-Event Simulator called YAFS [26].
+To simulate realistic Fog environments we used YAFS Discrete-Event Simulator [26].
 TABLE III: DDQL Hyper-parameters
@@
 Network Loss Function Huber Loss
-Flat interconnected Fog topologies are generated using the Networkx library [27] as random undirected graphs resembling the Internet Autonomous System network.
-Nodes in the graph are then identified as Cloud, Fog nodes, or APs based on their betweenness centrality [28].
-For instance, Fig.
-7 shows a sample topology with a Cloud, 11 Fog nodes, and 21 APs.
-Fog nodes are then assigned heterogeneous compute resources, in Instructions Per Second (IPS), from evenly spaced values in the range $[10^{3},10^{5}]$.
-This is done inversely proportional to the number of directly connected APs to resemble unbalanced load distribution.
-Figure 7: Unbalanced flat Fog topology with interconnected Fog nodes.
-The Cloud, with $10^{6}$ computing resources, is connected to two Fog nodes with the highest betweenness.
-Fog nodes are assumed closer to IoT devices with low link latency compared to the Cloud, which incurs higher delays due to longer transmission distances and additional processing layers.
-To simulate the physical distances of IoT-Fog, Fog-Fog, and Fog-Cloud links, the propagation delay of those links, in seconds, is drawn from a uniform distribution in the ranges $[1,2)$, $[2,4)$, and $[10,20)$, respectively [7].
-While their bandwidth, in Mbps, is drawn from a uniform distribution in the ranges $[10^{2},10^{3})$, $[10^{3},10^{4})$, and $[10^{3},10^{4})$, respectively [7].
-Heavy, moderate, and light workloads require $10^{4}$, $10^{3}$, and $10^{2}$ compute instructions, respectively.
-The data sizes of these workload categories are $10^{3}$, $10^{2}$, and $10^{1}$ Bytes, respectively.
-These values were chosen relative to Fog and link resources to create a system with scarce resources relative to the demand.
-Workloads are generated as a Poisson Point Process using exponential distributions to effectively model the stochastic nature of job arrivals in Fog computing environments [29].
-We used three scales for the exponential distributions, i.e., $\beta=$ 100, 150, and 200, to evaluate the performance under various load intensities, i.e., high, medium, and low generation rates/frequencies, respectively.
-Without efficient load distribution, the high rate can easily create a system bottleneck with heavy workloads and the unbalanced structure of the environment.
-In episodes of 10K simulation steps, training starts with the low rate for 30K training steps.
-A training step spans four decision steps, each spanning a variable number of simulation steps, depending on the workload generation rate.
-After 30K training steps, the rate instantly increases to medium, and then to high, requiring the agents to adapt to this rapid change using transfer learning.
-To evaluate the agents’ ability to reduce the number of queued jobs in each Fog node, we use the average system waiting delay as the main performance metric.
-We then evaluate balance and fairness by examining the similarity in waiting times and resource utilization across Fog nodes.
-Finally, we evaluate the system average for end-to-end workload execution delay, which includes network latency besides Fog waiting and service times.
-To underscore the benefits of learning independent policies over a common decision-making policy, we compare the proposed fully Distributed RL (DRL) approach against a Centralized single-agent RL (CRL) solution.
-We also include four baseline methods for completeness to show the significant difference between simple and intelligent load distribution solutions.
-These baselines include the selection of a random Fog node for each generated workload, and a Distributed Round-Robin (DRR) mechanism, where each AP assigns each workload among Fog nodes in a circular order.
-We also include the selection of the nearest Fog node to each workload’s AP.
-The final baseline is selecting the fastest Fog node for each workload based on compute requirements relative to node and link resources.
+Our distributed agents are evaluated on isolated Fog regions randomly generated with complex flat topologies, i.e., multiple interconnected Fog nodes.
+To realistically simulate Fog networks, each Fog region is created, using the Networkx library [27], from random undirected graphs resembling Internet Autonomous System networks.
+The nodes in these graphs are identified as the Cloud, Fog nodes, or IoT APs based on their betweenness centrality [28], i.e., the Cloud is in the center of the graph while APs are at the edges (see Fig.
+7).
+To resemble unbalanced load distribution relative to Fog resources, Fog nodes are assigned heterogeneous compute resources (uniformly between $[10^{3},10^{5}]$ Instructions Per Second (IPS)) inversely proportional to the number of their directly connected APs.
+Figure 7: A Fog region with unbalanced flat topology with 11 interconnected Fog nodes and 21 IoT APs.
+The Cloud is connected to two Fog nodes with the highest betweenness centrality and has computing resources of $10^{6}$ IPS.
+To simulate undedicated public networks with a distanced Cloud and Fog nodes towards the edge of the network, the propagation delays of IoT-Fog, Fog-Fog, and Fog-Cloud links are drawn from uniform distributions in the ranges $[1,2)$, $[2,4)$, and $[10,20)$ seconds, respectively [7].
+The bandwidths of those link categories are drawn from uniform distributions in the ranges $[10^{2},10^{3})$, $[10^{3},10^{4})$, and $[10^{3},10^{4})$ Mbps, respectively [7].
+To create a system with scarce resources, workloads with heavy, moderate, and light demands are simultaneously generated with compute requirements ($10^{4}$, $10^{3}$, and $10^{2}$ instructions, respectively) and data sizes ($10^{3}$, $10^{2}$, and $10^{1}$ Bytes, respectively) relative to Fog and link resources.
+These workloads are generated as a Poisson Point Process using exponential distributions to effectively model the stochastic nature of job arrivals in Fog computing environments [29].
+We used three scales for the exponential distributions, i.e., $\beta=$ 200, 150, and 100, to evaluate the performance under low, medium, and high load intensities, respectively.
+Without efficient load distribution, heavy workloads with high generation rates can easily create system bottlenecks due to the unbalanced Fog environment.
+This makes learning optimal load distribution strategies with high generation rates much harder than learning those strategies in environments with less demand.
+Therefore, our agents are trained with a low rate first for 30K training steps (in episodes of 10K simulation steps) before the rate instantly increases to the next load intensity.
+Using transfer learning, the agents rapidly adapt to the increased generation rates as they intensify from low to medium, and then from medium to high.
+In the following section, we compare the proposed fully Distributed RL (DRL) approach against a Centralized single-agent RL (CRL) solution to underscore the benefits of independent decision-making over learning a common policy.
+To demonstrate the effectiveness of intelligent load distribution compared to simple selection methods, we also included Random and Distributed Round-Robin (DRR) selection, as well as selecting the Nearest Fog node to each AP and the Fastest Fog service for each workload.
+The average system waiting delay is the performance metric that allows us to evaluate the ability to reduce workload accumulation in each Fog node.
+We also plot the variations in waiting times and resource utilization between all Fog nodes in the region, which helps evaluate the agents’ ability to balance the load in the network and provide fairness in resource utilization, respectively.
+We then measure the average end-to-end execution delay for the IoT application loop to analyze the impact of network latency and service times on load distribution decisions based solely on reducing waiting delay.
 V-A Performance Analysis
-The results presented in this section are averaged over 10 experiment trials with different random seeds to showcase the confidence intervals across different Fog topologies generated with the random graph generator.
-Using confidence intervals provides a measure of the variability in the results to assess the precision and reliability of the average values obtained from multiple experiment trials, enhancing the robustness of the findings.
-Due to the significant performance differences between the proposed DRL approach and the simple baseline methods, we used a logarithmic scale to show the values of DRL in the figures that compare it with baselines.
-The results are gathered using the highest generation rate, After training on the three rates each with 30K training steps, as it is the most difficult for the agent to learn [2].
-We use an evaluation episode of 100K simulation steps, i.e., 10×10$\times$10 × longer than training episodes to offer a more reliable measure of RL effectiveness and generalization capability in real-world scenarios.
-Starting with the average system waiting delay, i.e., the average waiting time in the queues of all Fog nodes, Fig.
-8LABEL:sub@fig:avgWaiting1 illustrates the performance of RL approaches compared to the baseline methods.
-In our evaluations, CRL performs better in reducing the system waiting delay than traditional baseline methods.
-This highlights the effectiveness of leveraging RL-based interactive learning to optimize load distribution in Fog environments.
-Furthermore, transitioning to DRL yields notable performance improvements beyond the CRL.
-By distributing decision-making across APs, the fully distributed independent agents capitalize on localized information and dynamic adjustments to the actions of other agents, resulting in even lower system waiting delays.
-(a) DRL compared to CRL and baseline methods
+The results in this section are averaged over 10 experimental trials, each with a different random seed, to illustrate the confidence intervals of performance across random variations in Fog architecture.
+These confidence intervals capture the variability in results due to topology changes, aiding in assessing our findings’ precision, reliability, and robustness across different environments.
+These results represent the performance in an evaluation episode that is 10×10$\times$10 × longer than training episodes, i.e., 100K simulation steps, offering a measure of generalization for lifelong inference during deployment.
+The workload generation rate in this evaluation episode is high to evaluate load distribution with frequently generated workloads, which is very difficult to learn [2].
+When compared against baseline methods, a logarithmic scale is used on the results to show the performance of RL-based solutions due to their significantly superior performance.
+Fig.
+8LABEL:sub@fig:avgWaiting1 illustrates the mean system waiting delay achieved using each load distribution method, which is the average time a workload must wait in the queue of a Fog node before being served.
+The results show superior performance with RL-based methods compared to simple baselines, highlighting the need for intelligent load distribution in unbalanced Fog environments.
+The results also show notable performance gains when transitioning from centralized to distributed RL, achieved by learning independent load distribution strategies that leverage resources in proximity differently for each distributed agent.
+Blindly learning separate policies by independent agents not only outperforms learning a common policy for all APs, but also eliminates the need to share network parameters, buffered interactions, or the actions of other agents.
+The small error bars in the figure show the robustness of RL-based solutions against variations in Fog topology across multiple experimental runs, showing the learning ability of the agents despite changes in problem complexity.
+(a) DRL vs. CRL and baselines
 (b) Realistic vs. unrealistic observations
@@
 Fig.
-8LABEL:sub@fig:avgWaiting2 focuses on the performance of both centralized and distributed RL solutions with respect to the frequency of observing the environment.
-It shows the performance of realistic observations using a Gossip protocol with 3 seconds interval between two consecutive actions.
-This is compared against the unrealistic assumption, commonly assumed in the literature, of readily available observations for every action.
-The figure highlights that realistic observations deliver slightly inferior performance compared to the optimal scenario where observations are unrealistically provided in real-time.
-This underscores the trade-off between realism and performance in RL-based load distribution solutions, where performance is directly related to the observation frequency.
-Despite the marginal performance degradation, realistic observation intervals align with practical constraints often encountered in real-world environments.
-After evaluating the average waiting delay in the system, we show the variations in waiting delay among all Fog nodes, where lower values represent similar waiting times across the Fog nodes.
+8LABEL:sub@fig:avgWaiting2 compares DRL with CRL only focusing on the impact of practical interval-based observations (observing Fog load every 3 seconds) compared to unrealistic real-time observations (a new observation for every action).
+This comparison is essential to show the trade-off between theoretical performance and realism of practical deployment constraints.
+This trade-off is often neglected in the literature by assuming Fog load information (from across the network/region) is readily available before every action.
+The figure shows that using the same observation to make multiple assignment decisions has a larger impact on DRL than CRL.
+The slightly inferior performance of DRL with interval-based observations, compared to the unrealistic optimal solution where observations are unrealistically provided in real-time, is due to the non-stationarity problem created by outdated information observed by independent agents.
+The impact on CRL with interval-based observations is smaller than DRL because the non-stationarity does not exist with CRL as the new observation depends on the behavior of a single agent.
 Fig.
-9LABEL:sub@fig:stdWaiting1 shows that RL-based solutions exhibit fewer variations than the baselines, with significantly better performance for DRL.
-By distributing RL decision-making, the agent in each AP manages the distribution of its workloads independently from other agents.
-This leads to a more effective load distribution considering more localized resources and adapting to the decisions of the other agents in the system.
-In contrast, CRL may exhibit higher variations in waiting time compared to DRL due to the centralized decision-making entity.
-The centralized agent struggles to efficiently adapt its decisions to the diverse workload demands from multiple APs across the heterogeneous Fog nodes.
-(a) DRL compared to CRL and baseline methods
+9LABEL:sub@fig:stdWaiting1 shows the variations in waiting delays between Fog nodes, measured using the standard deviation of waiting times across all Fog nodes.
+Lower values represent more equalized or balanced waiting times between Fog nodes.
+The error bars in the figure indicate the variability in these values across different experiments, demonstrating the level of result consistency under various Fog topologies.
+The results show how intelligent RL-based solutions can better balance the waiting times across Fog nodes than simple baselines, proving their ability to learn efficient load distribution strategies that minimize workload accumulation in all Fog nodes.
+By distributing RL decision-making, each agent can independently learn to distribute its load by leveraging nearby resources based on its location in the network.
+Therefore, DRL can balance the waiting times better than CRL since independently learning effective load distribution strategies, that fit the needs of each agent, is faster than learning a common policy for all APs.
+(a) DRL vs. CRL and baselines
 (b) Realistic vs. unrealistic observations
 Figure 9: The variations in the waiting delay between Fog nodes.
-Comparing DRL and CRL with respect to the frequency of observing the environment, Fig.
-9LABEL:sub@fig:stdWaiting2 shows a slightly different trend when it comes to the variations in waiting delay between Fog nodes.
-DRL with real-time observations is the best to equalize the waiting time across all Fog nodes.
-Using interval-based observations, DRL still performs better than CRL but with more variability across the different trials.
-CRL with real-time observations, however, shows the largest variability across trials, indicating less confidence,i.e., uncertainty, on the behavior of the agents during deployment.
-Interestingly, CRL with interval-based observations, with slightly less equalized waiting delay across Fog nodes, has less variability across trials than real-time observations.
-Fig.
-10LABEL:sub@fig:stdUtilization1 shows DRL’s ability to provide fair resource utilization between Fog nodes by reducing the variation in Fog resource utilization across the network.
-It shows significant improvement in DRL resource utilization fairness compared to CRL and other baselines.
-This mitigates the probability of overloading or underloading individual Fog nodes in the system.
-Maximizing the utilization of available resources enhances system performance by encouraging all Fog nodes to contribute equally to the overall workload processing, relative to their capacity constraints.
-On the other hand, mitigating overloading enhances system stability by ensuring the maximum availability of spare resources across the network, which minimizes the probability of dropping workloads in overloaded Fog nodes in case of a sudden surge in workload generation.
-(a) DRL compared to CRL and baseline methods
+To focus on the impact of interval-based observations on balancing the waiting times between Fog nodes, we again limit our comparison in Fig.
+9LABEL:sub@fig:stdWaiting2 to DRL and CRL only.
+The figure confirms that due to interval-based observations, outdated observations have a negligible impact on CRL compared to DRL.
+DRL is more affected by outdated observations because of the environment’s non-stationarity caused by the unobserved actions of multiple independent agents.
+With CRL, a single agent alters the state of the environment, making it stationary from that agent’s perspective, even if the same observation is used for multiple actions.
+However, DRL’s ability to balance waiting times between Fog nodes remains significantly better than CRL’s, with real-time and interval-based observations.
+With real-time observations, DRL demonstrates an impressive ability to achieve theoretically optimal load distribution, resulting in almost identical waiting times across all Fog nodes in the environment.
+Besides balancing waiting times between Fog nodes, RL-based solutions balance the consumption of Fog resources by reducing the variation in resource utilization between Fog nodes (see Fig.
+10LABEL:sub@fig:stdUtilization1).
+Balancing resource utilization provides fairness across Fog nodes and mitigates the probability of overloading or underloading individual Fog nodes in the region.
+This enhances system performance by encouraging all Fog nodes to equally contribute with their resources relative to their capacity and constraints.
+It also improves system stability by ensuring the maximum availability of spare Fog resources in the network, minimizing workload-dropping probability in case of a sudden surge in generation rates.
+While both RL methods are better than the baselines, DRL is significantly better in providing fair resource utilization using independently learned load distribution strategies.
+This proves the ability of our fully distributed agents to balance the system load without requiring the exchange of coordination information.
+(a) DRL vs. CRL and baselines
 (b) Realistic vs. unrealistic observations
 Figure 10: The variations in the module utilization between Fog nodes.
+Once again, we narrow down the comparison to DRL and CRL only to study the impact of interval-based observations on balancing resource utilization between Fog nodes (see Fig.
+10LABEL:sub@fig:stdUtilization2).
+The figure confirms our previous findings that the choice of observation frequency, i.e., real-time and interval-based observations, does not impact the performance of CRL due to having a single decision-making entity.
+Whether the CRL agent observes the environment before every action or makes an observation for all actions within 3 seconds, the variation in resource utilization (fairness) between Fog nodes is almost identical.
+DRL, however, is slightly impacted when observing the environment less frequently compared to unrealistically observing real-time Fog load before every action.
+Observing the load from Fog nodes once every 3 seconds is more practical for the actual implementation and deployment of the solution, and a slight degradation in the achievable performance can be tolerated given the constraints of real-world networks.
+To analyze the impact of network latency and service times on load distribution decisions based solely on reducing waiting delay, we need to show the average end-to-end execution delay for the whole IoT application loop.
+This will demonstrate the direct relationship between minimizing the waiting delay of individual workloads and minimizing the execution delay of the flow of workloads that make IoT applications.
 Fig.
-10LABEL:sub@fig:stdUtilization2 shows a small degradation in DRL fairness after using realistic interval-based observations.
-On the other hand, the choice between real-time and interval-based observations does not seem to have a significant difference in CRL fairness.
-Unlike Fig.
-9LABEL:sub@fig:stdWaiting2, Fig.
-10LABEL:sub@fig:stdUtilization2 shows small variability across the different trials, indicating less confidence on agents behavior.
-It is important to analyze agents’ ability to provide fairness in resource utilization, similar waiting delay between all Fog nodes, and minimal average waiting time.
-However, we still need to observe the average end-to-end execution delay, including network links latency, Fog queue waiting time, and Fog service time for all generated workloads.
-Fig.
-11LABEL:sub@fig:execution1 shows the ability of RL-based solutions to significantly minimize, compared to baseline methods, the average end-to-end execution delay for all generated workloads in the system.
-This performance improvement is directly attributed to agents’ ability to reduce the accumulation of workloads in the queues of Fog nodes, emphasizing the critical role of reducing the waiting delay in enhancing overall system performance.
-As shown earlier, there is a significant difference between DRL and CRL’s ability to minimize the variations in waiting delay and resource utilization between Fog nodes.
-However, the total execution delay achieved by both methods is quite similar, which is again attributed to their ability to reduce the waiting delay to a similar level relative to the performance of baseline methods.
-(a) DRL compared to CRL and baseline methods
+11LABEL:sub@fig:execution1 shows the ability of RL-based solutions to significantly minimize, compared to the simple baselines, the average execution time of the three application loops simultaneously running in the system.
+DRL performance is still slightly better than CRL due to the independent agents that learn different load distribution strategies based on their relative locations to Fog nodes, reducing workload accumulation in all Fog nodes better than using a single policy for the whole region.
+These independent agents find the optimal solution faster than learning a common policy by the single-agent solution.
+By avoiding the exchange of coordination information, and its associated overhead, these independent agents will also converge faster than implicitly learning a common policy using distributed agents that share their network parameters, experiences, or actions.
+(a) DRL vs. CRL and baselines
 (b) Realistic vs. unrealistic observations
 Figure 11: The average total execution delay.
-The relatively small difference in total execution delay between DRL and CRL (see Fig.
-11LABEL:sub@fig:execution2) underscores the efficacy of both approaches in optimizing system performance through different load distribution strategies.
-However, performance difference simply amplifies with less efficient solutions over longer simulations due to the accumulation of workloads over time, leading to increased total execution delays with DRL compared to CRL.
-Fig.
-11LABEL:sub@fig:execution2 also shows a small degradation in DRL performance using interval-based observations compared to CRL, but still performance better than CRL.
-The performance of CRL, on the other hand, is almost identical between real-time and interval-based observations.
-In summary, the results presented in this section offer insights into the performance of our proposed fully distributed MARL solution, i.e., DRL, against the performance of a single-agent RL solution and some other baselines.
-The results demonstrate the superior performance of RL approaches over baseline methods, with DRL showcasing the most notable improvements, highlighting the effectiveness of distributed decision-making in optimizing workload distribution.
-These improvements will amplify over longer simulations due to the accumulation of workloads over time, making such intelligent solutions the best option for real-world implementations.
-However, it’s crucial to note that the realistic performance of such solutions highly depends on observation frequency through multi-casting protocols, such as Gossip.
+With Fig.
+11LABEL:sub@fig:execution2, we can analyze the impact of interval-based observations on the average end-to-end execution delay of IoT application loops.
+Once again, the results show that real-time and interval-based observations lead to almost identical average execution delays using CRL.
+Even with outdated observations, the series of actions, by a single agent, between two consecutive observations determine the next observed state of the environment, leading to environment stationarity.
+Although DRL achieves a smaller average execution delay than CRL, DRL is slightly affected by these outdated observations due to having independent actors that simultaneously change the state of the environment.
+Even with interval-based observations, the performance of DRL is still better than CRL.
+The design of smaller regions can reduce DRL’s non-stationarity problem by minimizing the maximum communication delay for multi-casting Fog load information.
+This reduction enables the selection of smaller, yet practical, observation intervals based on the dimensionality of the region, lowering the impact of interval-based observations by independent decision makers.
+In summary, the results presented in this section demonstrate the superior performance of our proposed fully distributed MARL solution (DRL) compared to the CRL approach (decision-making using a common policy) and other baselines.
+The results highlight the ability of independently trained agents to learn distinct load distribution strategies more quickly than when learning a common policy.
+These strategies collectively reduce workload accumulation in Fog nodes more effectively than a CRL policy with the same amount of training.
+The fully distributed agents effectively minimize average waiting delay, balance waiting times, and ensure fair resource utilization across all Fog nodes in their region, thereby improving the average end-to-end execution delay.
+Without centralized or inter-agent coordination, our fully distributed agents can jointly learn optimal load distribution strategies more quickly and with less overhead than learning a common policy.
+A common policy is typically learned explicitly by a single agent or CTDE solutions, or implicitly through parameter, experience, or action sharing.
+Also, learning independent policies that optimize a common objective enables the management of smaller collaboration regions within a global network, while still ensuring global optimality.
+Using small collaboration regions allows for a scalable solution that reduces problem complexity since the agents have fewer Fog nodes to manage, i.e., state and action spaces are smaller, making learning easier and converging faster.
+The results also emphasize the trade-off between performance and real-world applicability, as we assess the impact of interval-based observations rather than assuming idealized conditions where Fog load information is continuously available before every action.
 VI Conclusion
-This paper presents a MARL fully distributed load-balancing solution that optimizes end-to-end execution time by minimizing Fog nodes’ waiting delay.
-The agents in our design work independently, and can dynamically adapt to environment changes using transfer learning in a life-long learning process.
-Our findings demonstrate the effectiveness of the proposed distributed solution in minimizing waiting time while enhancing resource utilization fairness across heterogeneous fog resources.
-By leveraging distributed decision-making, MARL agents outperform centralized solutions and traditional baselines, leading to significant improvements in overall system performance.
-Furthermore, we evaluate a more realistic approach to observing the state of the environment through an interval-based multi-casting protocol.
-Considering the real cost of observing the environment has never been considered before, and instead, observations are assumed to be unrealistically available in real-time for every decision to make.
-Unlike centralized load-balancing solutions, using fully distributed agents allows for a global-scale solution where groups of agents can work independently in small collaboration regions.
-For this reason, a primary focus for future research is to provide a real-world implementation to demonstrate the scalability and effectiveness of the proposed distributed load-balancing solution at a global scale.
-This can involve deploying independent agents in diverse Fog environments that span across distant geographical regions and heterogeneous network infrastructures.
-This research direction will bridge the gap between theoretical research and practical implementation, demonstrating the practical feasibility and effectiveness of the proposed solution at a global scale.
-In addition, testing with a real-world implementation can uncover potential limitations and constraints that can guide future research efforts in enhancing Fog load-balancing solutions.
+This paper presents a scalable Fog load balancing solution using fully distributed RL agents that independently learn to reduce workload accumulation in Fog nodes without exchanging learning or behavior information.
+Blindly optimizing the common objective of reducing Fog load, with independent strategies, allows our agents to reduce the average waiting delay, balance waiting times between Fog nodes, and provide fair resource utilization, leading to optimized end-to-end execution delay for IoT applications.
+Deploying an independent agent for each IoT AP leads to separate load distribution strategies, i.e., policies, that converge faster than learning a common policy for all APs.
+With limited training, our results show that learning independent policies outperforms explicitly learning a common policy by a single agent or CTDE-based distributed agents.
+Even with separate training, coordination between MARL agents implicitly leads to a common policy due to sharing network parameters, buffer experiences, or load distribution decisions.
+Unlike centralized and coordination-based load-balancing solutions, fully distributed agents allow global-scale implementations.
+The independent agents can be grouped to manage smaller sets of Fog nodes from the global network, forming Fog regions that might overlap by shared Fog resources.
+The ability of these agents to optimize a common objective in an isolated region without coordination guarantees their ability to optimize that objective even when some Fog resources are shared with another region.
+Managing fewer Fog nodes in each region reduces the state and action spaces, which makes learning easier and faster.
+To reduce required training further, the agents use a lifelong learning framework to adapt faster to environmental changes while mitigating the computational cost of continuous training.
+To do this, the agents are fine-tuned using transfer learning as generation rates significantly increase in the environment.
+Transferring policy network weights and replay buffer experiences between sources and target learning tasks leads to faster convergences and better generalization than learning from scratch.
+Each agent needs load information, i.e., for observations and rewards, from every Fog node in the network to make optimal load distribution decisions.
+Given communication latency, we can not assume the real-time availability of such information before every decision, especially with frequently generated workloads in large networks.
+Unfortunately, this assumption is common in the literature to report optimal performance that is not achievable in practice.
+Because the cost of collecting this information has never been considered before, we analyze a practical approach to observing Fog information in practical deployment.
+This analysis shows the trade-off between theoretical and practical performance by evaluating interval-based observations with realistic intervals from the Gossip multi-casting protocol.
+The findings demonstrate no impact on the single-agent solution and a slight deterioration in performance on the distributed agents.
+The unavoidable performance deterioration by independent agents is due to outdated observations, which make the environment non-stationary from the point of view of each agent, leading to harder learning.
+For future work, adaptive observation intervals could be explored as an alternative to fixed intervals, where the intervals are adjusted based on demand fluctuations.
+This approach could reduce overhead and mitigate the impact of outdated observations.
+Moreover, observation intervals could vary for each agent based on its location in the network relative to the Fog nodes it manages.
+Additionally, a real-world implementation is essential to validate the scalability and effectiveness of the proposed fully distributed load-balancing solution in practice.
+Deploying independent agents across geographically dispersed Fog environments with heterogeneous resources would provide a more accurate and realistic evaluation, especially in public networks with rapidly changing demands.
+This research direction will help bridge the gap between theoretical results and practical performance, demonstrating the feasibility of the proposed approach in real-world scenarios.